### v1

# Life-Code: Central Dogma Modeling with Multi-Omics Sequence Unification

Abstract
The interactions between DNA, RNA, and proteins are fundamental to biological processes, as illustrated by the central dogma of molecular biology. While modern biological pre-trained models have achieved great success in analyzing these macromolecules individually, their interconnected nature remains under-explored. In this paper, we follow the guidance of the central dogma to redesign both the data and model pipeline and offer a comprehensive framework, Life-Code, that spans different biological functions. As for data flow, we propose a unified pipeline to integrate multi-omics data by reverse-transcribing RNA and reverse-translating amino acids into nucleotide-based sequences. As for the model, we design a codon tokenizer and a hybrid long-sequence architecture to encode the interactions of both coding and non-coding regions with masked modeling pre-training. To model the translation and folding process with coding sequences, Life-Code learns protein structures of the corresponding amino acids by knowledge distillation from off-the-shelf protein language models. Such designs enable Life-Code to capture complex interactions within genetic sequences, providing a more comprehensive understanding of multi-omics with the central dogma. Extensive Experiments show that Life-Code achieves state-of-the-art performance on various tasks across three omics, highlighting its potential for advancing multi-omics analysis and interpretation.
Machine Learning, ICML
1 Introduction
The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins (Smith & Doe, 2023; Li & Wang, 2024; Li et al., 2024; Liu et al., 2024a). These endeavors aim to leverage deep learning to uncover functional elements, predict gene regulation, and accelerate protein engineering. Notably, the successful application of Transformer-based architectures in natural language processing (NLP) (Vaswani et al., 2017; Devlin et al., 2019) has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction (Zhou et al., 2022), protein structure elucidation (Jumper et al., 2021a), and more.
Table 1: Overview of Bio-Foundation Models. Unlike DNA models NT-2500M or EVO-7B using BERT or generative pre-training with pure DNA sequences, our Life-Code models the Central Dogma with the unified nucleotide input format. We consider four aspects of evaluations: The DNA tasks report the averaged MCC (%) over the Promoter Detection (all) and Covid Variants Classification, the RNA task considers the Splicing Site Reconstructed in the GUE benchmark, the Protein tasks report the averaged SRCC (%) Protein Fitness Prediction with DMS, and the Central Dogma evaluation considers the translation between CDS and Amino Acids with top-1 accuracy (%). Full results in Table A6.
Method NT-2500M EVO-7B LucaOne Life-Code
# Params 2500M 7B 110M 350M
Learning Tasks MLM MLM MLM+SL MLM+KD
Data Source DNA DNA Multi-omics DNA+AA
DNA Task 82.0 73.6 83.3 82.6
RNA Task 89.4 87.5 89.1 89.8
Protein Task 7.1 28.2 32.9 34.1
Central Dogma -- 75.5 84.8 85.6
Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM) severally proposed in each omics.
Figure 2: Illustration of Life-Code Framework, which contains three pipelines. (a) Data Pre-processing for the unified input with the central dogma. (b) Pre-training of the Tokenizer and the hybrid Encoder to model the contextual information and the central dogma rules. (c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning (SFT).
Despite the promise of these paradigms of foundation models in computational biology (Lin et al., 2023b; Nguyen et al., 2024a; Shen et al., 2024), three major obstacles continue to hamper multi-omics modeling: (1) Data Island. Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins (Hong et al., 2022), as shown in Figure 1. As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized. (2) Naïve “Copy and Paste”. While Transformers (Vaswani et al., 2017; Devlin et al., 2019) have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability (Rives et al., 2021). For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes. (3) Inefficiency for Long Sequences. Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length (AlQuraishi, 2019). This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
This paper introduces Life-Code, as shown in Figure 2, a unifying framework that addresses these limitations through multi-omics unification, bio-interpretable designs, and long-sequence efficiency. By grounding our method in the central dogma of molecular biology, we align DNA, RNA, and protein sequences under a single representation that captures cross-omics signals. We then augment this representation with a hybrid backbone architecture, where key biological structures (e.g., coding sequences, promoters, and regulatory motifs) guide the model to focus computational resources effectively. In particular, we explore specialized attention mechanisms that mitigate the memory bottleneck of naïve Transformer models while retaining sufficient global context for biologically relevant long-range dependencies. In summary, this work makes four key contributions to the field of computational biology with deep learning:
  • Multi-omics Modeling. We demonstrate a novel multi-omics approach by co-representing DNA, RNA, and protein within a single model, thus overcoming the traditional data island problem and enabling the transfer of knowledge across modalities.
  • Bio-interpretable Design. Our tokenization and training pipeline explicitly distinguishes coding (CDS) and non-coding (nCDS) regions, preserving biological interpretability. Through tasks like masked language modeling for nCDS and protein translation objectives for CDS, the model learns to capture both regulatory and translational signals.
  • Efficient Long-sequence Computing. We combine structured prior knowledge (e.g., double-helix complementarity, coding frames) with efficient linear attention mechanisms—balancing representational power with reduced complexity for long sequences. This hybrid strategy yields significant computational savings over naïve $\mathcal{O}(n^{2})$ Transformer approaches.
  • Comprehensive Validations. We evaluate Life-Code on diverse tasks in DNA, RNA, and protein applications. Empirical results confirm the effectiveness of our unified approach in these three fundamental modalities.
2 Method
2.1 Preliminaries
Central Dogma links islands.
The central dogma of molecular biology describes the flow of genetic information from DNA $\rightarrow$ RNA $\rightarrow$ Protein. DNA is transcribed into RNA, and RNA is subsequently translated into protein. This process ensures that the information encoded in the genome can be expressed to perform diverse biological functions. Let $\mathcal{D}\subseteq\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}^{*}$ be the space of DNA sequences, and let $\mathcal{R}\subseteq\{\texttt{A},\texttt{U},\texttt{C},\texttt{G}\}^{*}$ be the space of RNA sequences. For proteins, denote $\mathcal{P}\subseteq\Sigma^{*}$, where $\Sigma$ is the amino acid alphabet (e.g., $\Sigma=\{\texttt{Ala},\texttt{Arg},\texttt{Asn},...\}$). The central dogma stipulates two fundamental mappings:
$T_{\mathrm{transcribe}}:\mathcal{D}\to\mathcal{R},\quad T_{\mathrm{translate}}:\mathcal{R}\to\mathcal{P}.$ (1)
By composition, one obtains $T_{\mathrm{translate}}\circ T_{\mathrm{transcribe}}:\mathcal{D}\to\mathcal{P}$. In practice, DNA $\rightarrow$ RNA $\rightarrow$ Protein is not a strict one-to-one mapping due to splicing, alternative start sites, etc. However, these transformations offer critical biological correspondences that can link data from different omics sources. Why it matters: By grounding sequences in DNA ($\mathbf{x}\in\mathcal{D}$), we naturally align each transcript ($\mathbf{y}\in\mathcal{R}$) and its resultant protein ($\mathbf{z}\in\mathcal{P}$) within a unified coordinate system. This approach not only enlarges the effective dataset via cross-modality but also preserves mechanistic interpretability, as variations in $\mathbf{x}$ manifest as changes in $\mathbf{y}$ and $\mathbf{z}$.
The bi-directional hybrid model is more efficient.
DNA has three notable properties: (1) Length: A chromosome-scale DNA sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ can reach millions (or more) of base pairs (bp). (2) Double-Helix Complementarity: Each position $x_{i}\in\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}$ pairs with a complementary base on the opposite strand. (3) Distributed Regulatory Elements: Important functional sites (e.g., promoters, enhancers, splicing signals) can be located in different regions (CDS or nCDS). In summary, while Transformer-based architectures exhibit strong representational power, their attention mechanism scales with $\mathcal{O}(n^{2})$ complexity, making it difficult to handle sequences of length $n\gg 10^{4}$. Linear or sparse approximations (e.g., Mamba) reduce complexity but often sacrifice representational fidelity for complex biological contexts. Hence, the recent efficient hybrid strategy (Yang et al., 2025, 2024; Liu et al., 2024b, c) should be considered. We now formally describe how we unify DNA, RNA, and protein sources into nucleotides.
Figure 3: Data Sampling Pipeline. During pre-training, we sample the DNA sequences and the pair of CDS and amino acids from the DNA dataset and the DNA-AA paring dataset. Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample. As for the DNA sequence, we employ truncation and padding. As for the CDS and its corresponding amino acids, we apply the packing strategy (Warner et al., 2024) to sample with several CDSs.
2.2 Data Pipeline for Life-Code
Unified to nucleotides.
Given an RNA sequence $\mathbf{y}\in\mathcal{R}$ or a protein sequence $\mathbf{z}\in\mathcal{P}$, we map them back to DNA-level tokens. Specifically: For RNA, assume we have the corresponding genomic DNA region (e.g., from reference genome or aligned reads) and apply $\mathbf{y}=T_{\mathrm{transcribe}}(\mathbf{x})$ in reverse to identify the original $\mathbf{x}\in\mathcal{D}$. For protein, we leverage the standard codon translation table. Each amino acid $z_{j}\in\Sigma$ typically maps to one or more codons in $\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}^{3}$. We choose the canonical codon (or the most frequent codon) for each amino acid, yielding a surjective function $g:\Sigma\to\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}^{3}.$ Thus, a protein of length $L$ becomes a DNA-like sequence of length $3L$.
Data sampling.
Following GenBank (Benson et al., 2012) curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones. For CDS–Amino Acid pairs, we extract annotated CDS records and their corresponding protein sequences, then pack multiple short segments into a single training sample with special separators. See Figure 3 for a schematic illustration. This approach enables us to efficiently process both longer genomic segments and multiple shorter CDS-AA pairs in a unified framework. By balancing truncation/padding for DNA sequences and packing for CDS segments, we ensure that each batch contains a diverse mixture of long-range genomic context and coding information, thereby maximizing the model’s exposure to various biological signals.
Figure 4: Life-Code Tokenizer. To model the interactions among DNA, RNA, and Amino Acids, our tokenizer takes nucleotide acids (4 words) as the inputs, then takes codons as the latent vocabulary (64 words), and translates them to Amino Acids (20 words) as the outputs, which could be pre-trained by (a) nucleotide masked modeling and (b) CDS to Amino Acid translation.
Figure 5: Illustration of Life-Code Encoder pre-training pipeline. We model the long sequence by the hybrid encoder as a mixture of RNN blocks (DeltaNet) and self-attention blocks. With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma (with span masks), and (c) the knowledge distillation (KD) with pre-trained protein LM. Note that red lines indicates the learning objectives.
2.3 Life-Code Tokenizer Training
After mapping all sequences (DNA, RNA, and protein) to a unified DNA-like form, each sample can be viewed as a string in $\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}^{*}$. However, to effectively capture both nucleotide-level and codon-level patterns, as shown in Figure 4, our tokenizer incorporates a two-task training paradigm: (i) masked language modeling for nucleotide reconstruction and (ii) protein translation for CDS.
Tokenization with codon-level embeddings.
Consider a raw DNA segment of length $3L$. We first apply a 1-d convolution (kernel size $=3$) together with an “unfold” operation (stride $=3$) to transform each consecutive triplet $(x_{3j+1},x_{3j+2},x_{3j+3})$ into a codon-like representation:
$\mathbf{e}_{j}\;=\;\mathrm{Conv1d}(\,x_{3j+1:3j+3}\,).$
This step effectively condenses triplets—codons—into a more compact embedding sequence $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{L}\}\in\mathbb{R}^{L\times 3d}$. We apply a small linear projection or MLP to each $\mathbf{e}_{j}\in\mathbb{R}^{d}$, mapping it into the token embedding space.
Dual pre-training.
We jointly pre-train this tokenizer on two objectives: (1) Nucleotide MLM (Reconstruction). For each codon-embedded sequence, we randomly mask a subset of tokens and train the model to reconstruct the original nucleotides. Technically, we design a DNA De-Tokenizer that applies a Fold (stride = 3) and an inverse convolution (De-Conv1d) to recover the three-base codon, thereby reconstructing the full-length of DNA sequences. Formally, if $\widehat{\mathbf{x}}_{3j+1:3j+3}$ is the predicted codon for position $j$, the loss is:
$\mathcal{L}_{\mathrm{MLM}}\;=\;\sum_{j=1}^{L}-\log p\bigl{(}\mathbf{x}_{3j+1:3j+3}\,\big{|}\,\widetilde{\mathbf{x}}\bigr{)},$ (2)
where $\widetilde{\mathbf{x}}$ is the masked sequence. This ensures the tokenizer learns robust nucleotide representations that can reconstruct the original sequence. (2) CDS-to-Protein Translation. For coding regions, each codon corresponds to an amino acid ($\mathbf{z}_{j}\in\Sigma$). We thus include the Amino Acid Translator module, a two-layer MLP layer, that looks up the genetic code table and learns to predict $\mathbf{z}_{j}$ given $\mathbf{e}_{j}$:
$\mathcal{L}_{\mathrm{Trans}}\;=\;\sum_{j=1}^{L}-\log\,p\bigl{(}\mathbf{z}_{j}\,\big{|}\,\mathbf{e}_{j}\bigr{)}.$ (3)
By training on paired CDS–amino-acid data, the tokenizer acquires biologically grounded representations, effectively bridging codon-level features to protein-level semantics.
2.4 Bi-Directional Hybrid Model for Long Sequence
We describe the Life-Code Encoder, which takes the tokenized embeddings as input and models DNA sequences efficiently, respecting double-strand complementarity.
Bi-directional input.
To use the double-helix property as Schiff et al. (2024a), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
$(\mathbf{E}^{(+)},\ \mathbf{E}^{(-)})=\mathrm{Split}(\mathbf{E}),$
where $\mathbf{E}^{(+)}$ represents the forward strand embedding, and $\mathbf{E}^{(-)}$ represents the reverse (complementary) strand. We process these two strands in parallel using the same model weights or parameter-sharing scheme $f_{\theta}$:
$\mathbf{H}^{(+)}=f_{\theta}(\mathbf{E}^{(+)}),\quad\mathbf{H}^{(-)}=f_{\theta}(\mathrm{Reverse}(\mathbf{E}^{(-)})).$
Finally, we concatenate or fuse the representations:
[…]
yielding a representation that encodes both the forward and reverse complement features, where $\mathbf{H}\in\mathbb{R}^{D}$.
Hybrid Token Mixer.
Inspired by recent advances in EVO-style architectures (Nguyen et al., 2024a) and Gated Delta Networks (GDN) (Yang et al., 2024), we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA). Concretely, for every twelve encoder layers, eleven employ the GDN update rule—offering linear or near-linear complexity via kernel-based reordering—while one layer applies MHSA. The GDN layer relies on a delta update to maintain memory, reducing the quadratic overhead of naive Transformers. Meanwhile, the one full-attention layer provides additional global context, which is crucial for capturing fine-grained long-range dependencies. This 11 : 1 ratio of GDN to dense layers significantly boosts computational efficiency (compared to a pure $\mathcal{O}(n^{2})$ Transformer) while preserving model expressiveness and long-distance modeling capacity.
Figure 6: Empirical Analysis of Life-Code Tokenizer. Left: Codon Usage Bias in Life-Code Tokenizer across four representative species—E. coli, S. cerevisiae (yeast), D. melanogaster (fruit fly), and H. sapiens (human)—illustrating variations in codon frequency (%) for amino acids. The pastel strips highlight codons belonging to the same amino acid group. Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons (black crosses) form a distinct region. These views jointly demonstrate that Life-Code encodes both species-dependent codon usage preferences and biologically meaningful codon–amino acid relationships.
Encoder Pre-training.
Once the decoder and translator of our tokenizer weights are fixed (see Section 2.3), we train the Life-Code Encoder (Figure 5) on three tasks: (1) Masked DNA Reconstruction (MLM), (2) CDS-to-Amino-Acid Translation, and (3) Token-Level Knowledge Distillation (KD) from a pretrained Protein LM. The first two objectives match those used to pre-train the tokenizer but are now applied to the encoder output. Specifically, the encoder generates codon embeddings that a DNA De-tokenizer can reconstruct (MLM) and that an Amino Acid Translator can convert to protein residues (Translation).
For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues). Inspired by the DINO framework (Oquab et al., 2024), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (Lin et al., 2022). Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 2D\rightarrow D^{\prime}$. Then, the alignment or KD loss is defined as:
$\mathcal{L}_{\mathrm{KD}}\;=\;\sum_{j=1}^{L}-P_{\theta^{\prime}}\bigl{(}\mathbf{h}_{j}^{\mathrm{PLM}}\bigr{)}^{\top}\log P_{\theta}\bigl{(}\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}})\bigr{)}.$ (4)
We combine these three losses using a log-sum multi-task strategy (Lin et al., 2023a) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
$\log\bigl{(}\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr{)},$ (5)
where $\epsilon$ is a small constant. This adaptive weighting prevents any single task from dominating and ensures stable optimization across various training stages. Notably, we also adopt this strategy when training the tokenizer.
2.5 Empirical Consequences
Biological insights
In Figure 6, our analyses reveal that the Life-Code Tokenizer captures biologically meaningful codon relationships in two complementary ways. First, codon usage bias (left figure) indicates that different species, from E. coli to H. sapiens, vary markedly in their preference for certain codons—an effect traditionally tied to tRNA abundance and translational efficiency. By visualizing these preferences across amino acids, we observe how each organism clusters around specific high-frequency codons, underscoring the ecological and evolutionary factors shaping codon selection. Second, the t-SNE plot of the tokenizer’s learned codon embeddings (right figure) demonstrates that codons translating to amino acids with similar biochemical properties (e.g., hydrophobic or charged) are embedded close together, while the stop codons occupy a separate region. This structural arrangement confirms that the model internalizes codon–amino acid mappings, reflecting both frequency patterns and higher-level attributes like polarity or aromaticity. Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals—ranging from species-specific usage preferences to fundamental amino acid groupings.
Figure 7: Training Efficiency on a Single A100 GPU. We measure throughput (K t/s) at four sequence lengths–batch size settings. DNABERT-2 (blue) slows significantly at higher lengths and reaches an out-of-memory (OOM) error at $16\text{K}\times 2$. Both EVO (red) and Life-Code (green) scale more effectively, with Life-Code consistently achieving the highest throughput.
Table 2: Comparison on Genomic Benchmarks. Top-1 accuracy (%) averaged over several similar tasks is reported for popular DNA foundation models with SFT evaluation, where the best and the second best results are marked as the bold and underlined types.
Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
# Params (M) 6.6 86 117 113 498 7.9 93 100 1.7 350
Model Type RNN BERT BERT BERT BERT RNN BERT BERT CNN Hybrid
Enhancers (3 tasks) 80.88 80.14 82.81 83.22 84.56 79.96 82.37 82.79 80.90 85.27
Species Classification (2 tasks) 93.61 94.74 95.49 95.11 96.64 94.65 95.79 96.46 95.50 96.95
Regulatory Elements (3 tasks) 88.89 83.42 86.33 87.89 89.05 85.97 87.62 90.57 87.30 90.87
Table 3: Comparison on GUE Benchmark. MCC (%) averaged across several sub-tasks is reported with SFT evaluation for seven primary tasks, including Epigenetic Marks Prediction, Human Transcription Factor (TF) Prediction, Mouse Transcription Factor Prediction, Core Promoter Detection, Promoter Detection, Splice Site Reconstructed, and Covid Variants Classification (Virus Covid). The best and the second best results are marked as the bold and underlined types.
Method HyenaDNA DNABERT NT-2500M-multi DNABERT2 Caduceus-PS VQDNA MxDNA ConvNova Life-Code
# Params (M) 6.6 86 2537 117 1.9 93 100 1.7 350
Epigenetic Marks Prediction (10 tasks) 58.94 49.08 58.06 55.98 58.39 57.95 67.29 68.91 65.95
Human TF Detection (3 tasks) -- 64.17 63.34 70.11 -- 70.56 -- -- 71.26
Mouse TF Detection (3 tasks) -- 56.43 67.02 67.99 -- 69.80 -- -- 71.56
Core Promoter Detection (3 tasks) -- 71.81 71.63 70.53 -- 73.37 -- -- 73.51
Promoter Detection (3 tasks) -- 81.69 88.15 84.21 -- 86.58 -- -- 88.44
Splice Site Reconstructed -- 84.07 89.35 84.99 -- 89.53 -- -- 89.76
Virus Covid Classification -- 55.50 73.04 71.02 -- 74.32 -- -- 73.82
Training efficiency
We evaluated the throughput (thousands of tokens processed per second) of Life-Code, EVO, and DNABERT-2 on a single A100 GPU under increasing sequence lengths (with batch size adjusted accordingly). As shown in Figure 7, DNABERT-2 (blue) exhibits a steep drop-off at longer sequences and ultimately runs out of memory (OOM) at $16\text{K}\times 2$. In contrast, EVO (red) and Life-Code (green) both maintain higher throughput thanks to their hybrid attention mechanisms. Notably, Life-Code achieves the best scalability, retaining above 40K t/s even at $8\text{K}\times 4$. These results confirm that Life-Code’s efficient architecture can accommodate extensive input lengths with minimal degradation in training speed, making it well-suited for large-scale genomic modeling.
3 Experiments
We first introduce the network architectures and pre-training settings of our Life-Code tokenizer and encoder. Then, we evaluate Life-Code on DNA, RNA, protein, and multi-omics tasks with supervised fine-tuning (SFT) or zero-shot evaluation protocols. All experiments are conducted with PyTorch, transformers library, and NVIDIA A100-80G GPUs.
3.1 Experimental Setup
Life-Code Tokenizer.
With the nucleotide vocabulary, the Life-Code tokenizer contains the following modules: (i) A linear projection from the DNA input to 384-dim implemented by nn.Embedding; (ii) A GDN block (DeltaNet) with 384-dim for global contextual modeling; (iii) A 1-d Convolution with a kernel size of 3 and stride of 1, followed by an UnFold operation to merge every three nucleotide tokens into 768-dim codon embedding. Similarly, we also design the DNA De-Tokenizer and Amino Acid Translator for tokenizer pre-training. As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (Loshchilov & Hutter, 2019) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2. As shown in Table A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases as described in Figure A1 and Appendix A. View Appendix A.2 for details.
Life-Code Encoder.
Similar to the BERT-Large architecture (Devlin et al., 2019), the Life-Code encoder has 24 layers in total with an embedding dim of 1024 and 340M parameters. We enhance the model with three-fold designs: (i) Mixture of DeltaNet (Yang et al., 2025) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block. (ii) The LLaMA-like macro design (Touvron et al., 2023) with RMSNorm (Zhang & Sennrich, 2019), Layer Scale (Touvron et al., 2021), Rotary Position Embedding (RoPE) (Su et al., 2021), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences. (iii) During pre-training, we apply the packing strategy (Warner et al., 2024) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences. With three pre-training tasks in Eq. 5, the Life-Code tokenizer and encoder are optimized by AdamW optimizer for 1M iterations with the batch size of 256 and the basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for KD from a pre-trained Protein LM, we adopted ESM2-650M (Lin et al., 2022) and a protein decoder with an output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler. View Appendix A.2 for more details.
3.2 Comparison on Genomic Benchmarks
As DNA sequences are the foundation of the central dogma, we primarily compare with latest DNA foundation models with the SFT evaluation protocol proposed by (Nguyen et al., 2024c; Zhou et al., 2024), following benchmark evaluations for fair comparison. We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (Ji et al., 2021a), DNABERT2 (Zhou et al., 2024), GENA-LM (Fishman et al., 2023), Nucleotide Transformer (NT) (Dalla-Torre et al., 2023), with different tokenizers like VQDNA (Li et al., 2024) and MxDNA (Qiao et al., 2024); (b) RNN-based linear attentions including HyenaDNA (Nguyen et al., 2024c) and Caduceus (Schiff et al., 2024a); (c) CNN model ConvNova (Anonymous, 2025).
Genomic Benchmarks.
We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (Grešová et al., 2023), which we adopted results from GenBench implementations (Liu et al., 2024a) and all compared DNA models are fully SFT and reported the top-1 accuracy. Table 2 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages of both RNN and self-attention blocks.
GUE Benchmarks.
We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (Zhou et al., 2024), as shown in Table 3, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification. All results are adopted from DNABERT2 or each paper with their official SFT setups, and Matthews Correlation Coefficient (MCC) is reported as the metric. Compared to popular DNA models, our Life-Code shows superiority in gene annotation tasks, which are important regulation processes in the central dogma, while achieving competitive results in Epigenetic regulating tasks.
Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. Top-1 AUC score or F1 score are reported for DNA models or RNA models with SFT evaluation, respectively. The best and the second best results are marked as the bold and underlined types.
Dataset SpliceAI DNABERT2 NT Caduceus Life-Code
SpliceAI 63.2 67.1 63.9 69.1 69.6
Dataset Spliceator DDSP RNA-FM RINALMo Life-Code
Spliceator 91.0 91.9 92.1 94.8 94.5
3.3 Comparison on RNA Splicing Site Prediction
Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks. We consider two typical datasets with the SFT evaluation protocol. As for SpliceAI dataset (Jaganathan et al., 2019), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma. As for Spliceator dataset (Scalzitti et al., 2021), we then consider RNA splicing methods (Spliceator, DDSP (Naito, 2019), RINALMo (Penic et al., 2024)) and RNA language model, RNA-FM (Shen et al., 2024), with RNA data of multi-species. As shown in Table 4, our Life-Code achieves competitive F1 scores with the specially designed splicing method and surpasses RNA-FM.
Table 5: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets. The best and the second best results are marked as the bold and underlined types. ESM2 is marked as gray since it takes the amino acids as the input, while other models utilize the gene sequences.
Method GenSLM NT-2500M EVO ESM2-650M Life-Code
# Params 2.5B 2.5B 7B 650M 350M
Bacteria 24.7 9.4 45.3 51.2 45.7
Human 6.9 4.7 11.1 37.5 22.4
3.4 Comparison on Protein Fitness Prediction
Following EVO (Nguyen et al., 2024a), we then evaluate foundation models pre-trained with DNA or protein data with Protein Fitness Prediction tasks using zero-shot evaluation protocol (without task-specific SFT). Deep Mutational Scanning (DMS) datasets (Notin et al., 2022; Nguyen et al., 2024a) introduce a comprehensive set of mutations into protein CDS to measure their effects on fitness, where fitness serves as a metric for how effectively a protein performs a specific function. As shown in Table 5, bacterial (prokaryote) and human (eukaryote) protein fitness is evaluated by Spearman’s Rank Correlation Coefficient (SRCC), where the protein specialist model ESM2 achieves the best result. Life-Code outperforms other DNA models like GenSLM (Zvyagin et al., 2022) and EVO-7B with fewer parameters, which might be attributed to its modeling of the translation and folding processes in the central dogma.
Table 6: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation. Bold type denotes the best result.
Method ncRNA-Protein Central Dogma
DNABERT2+ESM2 93.32 73.31
EVO -- 75.45
LucaOne 93.80 84.84
Life-Code 93.45 85.63
3.5 Comparison on Multi-omics Tasks
We conduct further comparisons with proposed biological foundation models (EVO (Nguyen et al., 2024a) and LucaOne (He et al., 2024)) with multi-omics tasks. Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins. As shown in Table 6, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve the competitive result as the supervised LucaOne, which should process two inputs with separate modules. Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and the protein sequences, where Life-Code achieves the best accuracy over existing models.
Table 7: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory on the single A100-80G.
Method # Params 1k 20k 32k 100k 250k
HyenaDNA 6.6M 61.1 87.4 93.4 96.4 97.9
DNABERT2 117M 61.0 86.8 99.3 OOM OOM
Life-Code 350M 62.2 87.6 99.3 99.4 99.5
3.6 Empirical Analysis
After verifying the effectiveness with extensive comparison results, we further consider long-sequence tasks like HyenaDNA (Nguyen et al., 2024c) in addition to Figure 7. We compare the inference throughput of our bi-directional hybrid model to baselines (FlashAttention and HyenaDNA). In Table 7, empirical profiling shows that LifeCode achieves competitive speed with these specialized methods while retaining strong global modeling capacity. In particular, mixing a small fraction of MHSA enables better downstream performance without incurring a large computational penalty, e.g., splicing site prediction in Table 4.
Table 8: Ablation Study of Life-Code tokenizer, encoder, and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
Tokenizer Encoder PT Tasks DNA Protein CD
Linear (1-mer) MLM 58.3 -- 28.7
GDN (codon) MLM+Trans. 84.5 6.4 63.2
GDN (codon) GDN MLM+Trans. 95.0 13.8 73.9
GDN (codon) Hybrid MLM+Trans. 97.2 15.4 75.7
GDN (codon) Hybrid MLM+Trans.+KD 97.0 34.1 85.6
3.7 Ablation Study
We conduct ablation on the pre-training tasks and the network designs for the tokenizer and encoder, as shown in Table 8. We consider the top-1 accuracy (%) Covid Variants Classification in the GUE benchmark for the DNA task, the averaged SRCC (%) of bacterial and human protein fitness with DMS used in EVO as the protein task, and the top-1 accuracy (%) of the central dogma evaluation task proposed by LucaOne. Firstly, we verify that the codon tokenizer (instead of 1-mer) could learn more global contextual information by the RNN module and Conv1D (with a kernel size of 3) pre-trained by the MLM task and the cDNA-AA translation task (Trans.). Then, we verified the effectiveness of the hybrid encoder and found that knowledge distillation from ESM2-650M could benefit the protein-related tasks.
4 Related Works
DNA Foundation Models.
Recent years have seen rapid progress in DNA foundation models that leverage Transformer-like architectures. Early attempts such as DNABERT (Ji et al., 2021b) treated genomic sequences akin to linguistic tokens, enabling pre-trained contextual representations. Follow-up work DNABERT-2 (Zhou et al., 2023) expanded this paradigm by introducing more efficient training protocols and supporting multi-species data. In parallel, The Nucleotide Transformer (Dalla-Torre et al., 2023) demonstrated the feasibility of scaling up Transformer architectures for human genomics. Beyond Transformers, kernel-based or hierarchical approaches emerged—e.g., HyenaDNA (Nguyen et al., 2024b), which reduces the quadratic complexity of attention for extremely long sequences. Meanwhile, alternative designs like Caduceus (Schiff et al., 2024b) incorporate selective structural priors for long-range DNA modeling, and MxDNA (Qiao et al., 2024) explores adaptive tokenization schemes that automatically discover suitable patterns for genomic data.
Protein and RNA Advances.
On the protein side, AlphaFold series (Jumper et al., 2021a, b) revolutionized structure prediction, catalyzing a surge in protein-based language models such as ESM-2 (Lin et al., 2022), which refines large-scale protein embeddings. Beyond proteins, RNA modeling also gained traction; for instance, (Shen et al., 2024) employed a language-model-based technique for accurate 3D structure prediction. These efforts underscore the trend toward specialized architectures for each biomolecule yet also highlight the desire for integrative multi-omics solutions.
Multi-Omics Modeling.
Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks. CD-GPT (Zhu et al., 2024) explicitly connects DNA, RNA, and proteins through the central dogma, while BSM (Xiang et al., 2024) highlights the potential for small but effective models covering genes and proteins simultaneously. Evo (Nguyen et al., 2024a) similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach. Closely related is LucaOne (He et al., 2024), which advocates a unified nucleic acid and protein language for biological representation.
5 Conclusion and Limitations
We have introduced Life-Code, a method that unifies DNA, RNA, and protein sequences by mapping the latter two back to a DNA-like representation. This design offers straightforward multi-omics integration and maintains biological interpretability via differential encoding of coding versus non-coding regions. To handle the long-range dependencies inherent in genomic data, we employ an efficient symmetric convolution-based architecture. Moreover, knowledge distillation from large protein models allows our approach to scale without excessive resource demands.
Limitations
Experimental results demonstrate that Life-Code achieves competitive performance in protein structure prediction and phenotype analysis. However, limitations remain: refining the modeling of heterogeneous non-coding regions, incorporating post-translational modifications, and addressing the loss of fine-grained information during knowledge distillation. Future work will also focus on integrating additional omics layers (e.g., epigenomics, metabolomics) for a deeper understanding of biological systems. Despite these challenges, we believe Life-Code offers a promising step toward comprehensive multi-omics data.
Impact Statements
This paper presents a novel approach, Life-Code, aimed at advancing the field of machine learning and bioinformatics by unifying the modeling of multi-omics data through the guidance of the Central Dogma of molecular biology. By integrating DNA, RNA, and protein data into a single framework, our work has the potential to improve the understanding of complex biological processes and drive advancements in genomics, proteomics, and systems biology.
Ethical Aspects
This work has significant potential applications in fields such as personalized medicine, drug discovery, and disease diagnostics. However, we acknowledge the need to address concerns related to data privacy and the responsible use of sensitive genetic information. We have adhered to ethical guidelines in our experiments, ensuring that no private or identifiable data was used.
Appendix
  • In Appendix A, we provide implementation details of pre-training datasets, network architectures, and training schemes of pre-training and fine-tuning stages with hyper-parameter settings.
  • In Appendix B, we provide detailed descriptions of downstream tasks of biological applications and full comparison results.
Figure A1: Data Collection Pipeline. We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA paring datasets.
Appendix A Implementation Details
Table A1: Configuration of pre-training datasets. As for the DNA dataset, we collect DNA/RNA sequences from the NCBI RefSeq database. As for the DNA-AA pairing dataset, we collect the cDNA (coding sequences) and its corresponding Amino Acids (or using translation) from the GenBank database and also collect Amino Acids and their corresponding cDNA sequences using reverse translation from the UniRef50 database.
Dateset Data Type Seq Count Data Source
DNA DNA 51,257,875 NCBI RefSeq
RNA 6,463,852 NCBI RefSeq
cDNA-AA pair 16,786,593 GenBank
DNA-AA AA-cDNA pair 17,245,138 UniRef50
mRNA-AA pair 3.908,074 GenBank
A.1 Pre-training Dataset
We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1. As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer (Dalla-Torre et al., 2023) and DNABERT2 (Zhou et al., 2024). As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma. We also collect some Amino Acids in the UniRef50 database following LucaOne (He et al., 2024) and obtain their corresponding cDNA by reverse translation with online tools. We provide detailed information for used datasets in Table A1.
A.2 Life-Code Tokenizer
Vocabulary.
There are two vocabularies used in Life-Code. The unified vocabulary only uses 4 nucleotides {A, T/U, C, G} of nucleic acid with 5 special tokens, including “[U]”/“[UNK]”, “[PAD]”, “[CLS]”, “[SEP]”, and “[MASK]” for unknown nucleotides, padding tokens, the class token, separator tokens, and masking tokens. Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”). It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS). The pre-trained protein language model employs the amino acid vocabulary of ESM-2 (Lin et al., 2022).
Table A2: Configuration of the network designs and pre-training settings for Life-Code models. GDN blocks denote the Gated DeltaNet block with linear complexity, while Attention blocks denote the self-attention block with FLASH-Attention implementation.
Configuration Tokenizer Encoder
Embedding dim 384 1024
Block number 1 24
GDN blocks 1 22
Attention blocks 0 2
Attention heads 0 24
Parameters 8M 340M
Optimizer AdamW
$(\beta_{1},\beta_{2})$ $(0.9,0.98)$
Training iterations 100,000 1,000,000
Weight decay $1\times 10^{-2}$
Base learning rate $2\times 10^{-4}$ $1\times 10^{-4}$
Batch size 512 256
LR scheduler Cosine Annealing
Warmup iterations 5000 10,000
Gradient clipping 1.0
Tokenizer Network.
As shown in Figure 4, with the nucleotide vocabulary (including 4 nucleic acids and 5 special symbols), the Life-Code tokenizer contains the following modules: (a) A linear projection from 9-dim to 384-dim implemented by nn.Embedding; (b) A GDN block (Gated DeltaNet) with 384-dim for global contextual modeling with linear computational complexity; (c) A 1-d Convolution with a kernel size of 3 and stride of 1, followed by an UnFold operation to merge every three nucleotide tokens into 768-dim codon embedding. Similarly, we design the DNA De-Tokenizer with the symmetrical network as the Life-Code tokenizer: (a) Fold operation with a 1-d Convolution with a kernel size of 3 to unmerge the codon embedding to 384-dim, (b) A linear projection from 384-dim to 9-dim vocabulary to reconstruct the original DNA sequences. We also design the Amino Acid Translator as a two-layer MLP that translates the codon embedding to the corresponding Amino Acid sequences.
Pre-training Settings.
As shown in Figure 4, we pre-train the Life-Code tokenizer with the DNA de-tokenizer and Amino Acid de-tokenizer by AdamW optimizer for 100,000 iterations (randomly sampled datasets) with a basic learning rate of 2e-4 and a batch size of 512, as detailed in Table A2. We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 8 and a gradient accumulation time of 4.
A.3 Life-Code Encoder
Encoder Architecture.
As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet (Yang et al., 2025)) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 (MiniMax et al., 2025), which could utilize the complementary properties of GDN and MHSA while maintaining efficiency. (2) The model macro design employs pre-norm (Wang et al., 2019a) with RMSNorm (Zhang & Sennrich, 2019), Layer Scale (Touvron et al., 2021), Rotary Position Embedding (RoPE) (Su et al., 2021), SwiGLU (Touvron et al., 2023), and FlashAttention implementations to facilitate training large-scale models stably with long sequences. (3) During pre-training, we apply the packing strategy (Warner et al., 2024) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
Pre-training Settings.
As shown in Figure 5, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq. 5 for 1M steps with the batch size of 256 and the basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D(Lin et al., 2022) and a protein decoder with the output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler (decay to $1\times 10^{-6}$). We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 2 and a gradient accumulation time of 16.
A.4 Supervised Fine-tuning
In most cases, we apply Supervised Fine-tuning (SFT) to transfer pre-trained models to downstream tasks. Following (Nguyen et al., 2024c; Zhou et al., 2024), adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy (Hu et al., 2021) is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32. For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate {1e-5, 5e-5, 1e-4}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench (Liu et al., 2024a). Note that the maximum input length will be determined for different tasks since the sequence lengths of downstream tasks vary widely. We report the averaged results over three runs with the optimal settings.
Table A3: Full Results on Genomic Benchmarks. Top-1 accuracy (%) averaged across three trials is reported for the latest DNA foundation models, where the best and the second best results are marked as the bold and underlined types.
Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
# Params (M) 6.6 86 117 113 498 7.9 93 100 1.7 350
Mouse Enhancers 79.34 80.99 81.82 82.97 85.12 81.63 81.06 80.57 78.40 85.46
Human Enhancers Cohn 72.96 70.23 75.87 75.63 76.12 73.76 75.63 74.67 74.30 76.85
Human Enhancers Ensembl 90.33 89.19 90.75 91.07 92.44 84.48 90.41 93.13 90.00 93.49
Coding vs Intergenomic 90.97 93.64 93.58 93.24 95.76 93.72 94.35 95.28 94.30 96.14
Human vs Worm 96.24 95.84 97.39 96.98 97.51 95.57 97.23 97.64 96.70 97.75
Human Regulatory 93.08 88.16 87.94 88.10 93.79 87.30 90.92 94.11 87.30 93.93
Human OCR Ensembl 79.14 74.96 75.82 78.98 80.42 81.76 76.58 81.05 79.30 82.02
Human NonTATA Promoters 94.45 87.13 95.24 96.60 92.95 88.85 95.37 96.56 95.30 96.65
Table A4: Full Results on GUE benchmark. MCC (%) is reported for Epigenetic Marks Prediction, Human Transcription Factor (TF) Prediction, Mouse Transcription Factor Prediction, Core Promoter Detection, Promoter Detection, Splice Site Reconstructed, and Covid Variants Classification (Virus Covid). The best and the second best results are marked as the bold and underlined types.
Method HyenaDNA DNABERT NT-2500M-multi DNABERT2 Caduceus-PS VQDNA MxDNA ConvNova Life-Code
# Params (M) 6.6 86 2537 117 1.9 93 100 1.7 350
Human TF-0 -- 66.84 66.64 71.99 -- 72.48 -- -- 71.58
Human TF-1 -- 70.14 70.28 76.06 -- 76.43 -- -- 75.92
Human TF-2 -- 61.03 58.72 66.52 -- 66.85 -- -- 70.63
Human TF-3 -- 51.89 51.65 58.54 -- 58.92 -- -- 58.74
Human TF-4 -- 70.97 69.43 77.43 -- 78.10 -- -- 79.45
Mouse TF-0 -- 44.42 63.31 56.76 -- 58.34 -- -- 64.10
Mouse TF-1 -- 78.94 83.76 84.77 -- 85.81 -- -- 86.51
Mouse TF-2 -- 71.44 71.52 79.32 -- 80.39 -- -- 80.49
Mouse TF-3 -- 44.89 69.44 66.47 -- 69.72 -- -- 71.25
Mouse TF-4 -- 42.48 47.07 52.66 -- 54.73 -- -- 55.46
Core Promoter (all) -- 68.90 70.33 69.37 -- 71.02 -- -- 70.69
Core Promoter (no TATA) -- 70.47 71.58 68.04 -- 70.58 -- -- 71.05
Core Promoter (TATA) -- 76.06 72.97 74.17 -- 78.50 -- -- 78.78
Promoter (all) -- 90.48 91.01 86.77 -- 90.75 -- -- 91.33
Promoter (no TATA) -- 93.05 94.00 94.27 -- 94.48 -- -- 95.03
Promoter (TATA) -- 61.56 79.43 71.59 -- 74.52 -- -- 78.97
Splice Reconstructed -- 84.07 89.35 84.99 -- 89.53 -- -- 89.76
H3 78.14 73.10 78.77 78.27 77.90 79.21 82.14 81.50 81.28
H3K14ac 56.71 40.06 56.20 52.57 54.10 54.46 68.29 70.71 68.41
H3K36me3 59.92 47.25 61.99 56.88 60.90 61.75 65.46 68.31 67.25
H3K4me1 44.52 41.44 55.30 50.52 48.80 53.28 54.97 56.60 57.32
H3K4me2 42.68 32.27 36.49 31.13 38.80 34.05 55.30 57.45 50.31
H3K4me3 50.41 27.81 40.34 36.27 44.00 39.10 63.82 67.15 53.97
H3K79me3 66.25 61.17 64.70 67.39 67.60 68.47 73.74 72.08 72.26
H3K9ac 58.50 51.22 56.01 55.63 60.40 56.63 63.15 68.10 65.45
H4 78.15 79.26 81.67 80.71 78.90 81.84 80.89 81.12 81.89
H4ac 54.15 37.24 49.13 50.43 52.50 50.69 65.14 66.10 61.37
Virus Covid Classification -- 55.50 73.04 71.02 -- 74.32 -- -- 73.82
Appendix B Downstream Task Settings and Extensive Comparison Results
B.1 DNA Tasks with Genomics Benchmark
As proposed by (Grešová et al., 2023), three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark. As for the enhancer prediction, three datasets are provided for identifying enhancer regions in the mouse or human genome. As for the species classification, two datasets are selected for identifying sequences as either coding (exonic) or intergenic (non-coding) and classifying sequences as originating from humans or worms (C. elegans). As for the regulatory elements classification, three datasets are used for classifying sequences as regulatory regions based on Ensembl annotations, identifying open chromatin regions, or identifying non-TATA promoter regions in the human genome. We utilize the fully reproduced results of various DNA models in GenBench (Liu et al., 2024a).
B.2 DNA Tasks with GUE Benchmark
As proposed by DNABERT2 (Zhou et al., 2024), the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric. To comprehensively evaluate the genome foundation models in modeling variable-length sequences, tasks with input lengths ranging from 70 to 1000 are selected. The following descriptions of the supported tasks are included in the GUE benchmark, where these resources are attached for illustration.
Promoter Detection (Human). This task identifies human proximal promoter regions essential for transcription initiation. Accurate detection aids in understanding gene regulation and disease mechanisms. The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) (Dreos et al., 2013). Meanwhile, we construct the non-promoter class with equal-sized randomly selected sequences outside of promoter regions but with TATA motif (TATA non-promoters) or randomly substituted sequences (non-TATA, non-promoters). We also combine the TATA and non-TATA datasets to obtain a combined dataset named all.
Core Promoter Detection (Human). This task is similar to the detection of the proximal promoter with a focus on predicting only the core promoter region, the central region closest to the TSS, and the start codon. A much shorter context window (center -34 +35 bp around TSS) is provided, making this a more challenging task than the prediction of the proximal promoter.
Transcription Factor Binding Site Prediction (Human). This task predicts human transcription factor binding sites (TFBS), crucial for gene expression regulation. Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) (Consortium et al., 2012b) are collected via the UCSC genome browser. TFBS sequences are 101-bp regions around peaks, while non-TFBS sequences match in length and GC content. There are 5 datasets selected from a curated subset of 690, excluding trivial or overly challenging tasks.
Splice Site Prediction (Human). This task predicts splice donor and acceptor sites, the exact locations in the human genome where alternative splicing occurs. This prediction is crucial to understanding protein diversity and the implications of aberrant splicing in genetic disorders. The dataset (Scalzitti et al., 2021) consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome. As suggested by Ji et al. (2021a), existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality. As such, we reconstruct the dataset by iteratively adding adversarial examples (unseen false positive predictions in the hold-out set) in order to make this task more challenging.
Transcription Factor Binding Site Prediction (Mouse). This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) (Consortium et al., 2012a) from the UCSC genome browser. Negative examples are created by di-nucleotide shuffling. Five datasets are randomly selected from the 78 datasets using the same process as the human TFBS prediction dataset.
Epigenetic Marks Prediction (Yeast). This task predicts epigenetic marks in yeast, which influence gene expression without altering DNA sequences. Precise prediction of these marks aids in elucidating the role of epigenetics in yeast. We download the 10 datasets from http://www.jaist.ac.jp/~tran/nucleosome/members.htm and randomly split each dataset into training, validation, and test sets.
Covid Variant Prediction (Virus). This task aims to predict the variant type of the SARS_CoV_2 virus based on 1000-length genome sequences. We download the genomes from the EpiCoV database (Khare et al., 2021) of the Global Initiative on Sharing Avian Influenza Data (GISAID). We consider 9 types of SARS_CoV_2 variants, including Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda and Zeta.
Table A5: Full Results of mRNA Splicing Site Prediction. With two splicing datasets proposed by SpliceAI and Spliceator, the top-1 AUC score or F1 score is reported with DNA models or RNA models, respectively. The best and the second best results are marked as the bold and underlined types.
Dataset SpliceAI DNABERT2 NT GENA-LM Caduceus Life-Code
Donor 57.4 63.5 55.7 62.9 64.2 64.3
Acceptor 69.1 70.7 72.2 73.0 74.0 74.6
Mean 63.2 67.1 63.9 67.9 69.1 70.0
Dataset Spliceator SpliceFinder DDSP RNA-FM RINALMo Life-Code
Human 90.0 84.5 90.6 90.7 91.3 91.5
Fish 91.9 91.8 93.6 93.7 97.4 97.3
Fly 91.0 84.2 91.4 91.9 95.8 94.7
B.3 mRNA Splicing Tasks
Following (Zhou et al., 2024; Shen et al., 2024), we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression. During splicing, introns are removed from precursor messenger RNAs (pre-mRNAs), and exons are joined together to form mature mRNAs. This process is essential for generating functional mRNAs that could be translated into proteins. Identifying splice sites—the donor sites at the 5’ end of introns and the acceptor sites at the 3’ end—is vital for accurately predicting gene structure and location. Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset (Wang et al., 2019b) used in the GUE benchmark.
Spliceator dataset. This dataset (Scalzitti et al., 2021) consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans. The gold standard dataset GS 1 is adopted, which contains an equal number of positive and negative samples, and the F1 score is used as the evaluation metric. We chose three independent test datasets containing the samples from 3 different species of humans, fish (Danio rerio), and fruit fly (Drosophila melanogaster).
SpliceAI dataset. This dataset (Jaganathan et al., 2019) also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information. SpliceAI also references the ClinVar database to evaluate the clinical significance of predicted splicing variants, which contains information on clinically relevant variants and their associations with diseases. This dataset can be regarded as a long-range evaluation and adopts the top-1 AUC-ROC score as the metric.
Table A6: Full Results of Foundation Models. As for DNA and RNA tasks, MCC (%) is reported for Promoter Detection (all), Covid Variants Classification, and Splice Site Reconstructed. SRCC (%) is reported for Bacterial and Human Protein Fitness Prediction with DMS. Top-1 accuracy (%) is reported for the Central Dogma evaluation. The best result is marked as the bold type.
Method NT-2500M EVO-7B LucaOne Life-Code
Promoter (all) 91.0 88.5 91.6 91.3
Virus Covid Classification 73.0 58.7 75.1 73.8
Splice Reconstructed 89.4 87.5 89.1 89.8
Bacterial Protein 9.4 45.3 46.1 45.7
Human Protein 4.7 11.1 19.6 22.4
Central Dogma -- 75.5 84.8 85.6
B.4 Protein and Multi-omic Tasks
Zero-shot Protein Fitness Prediction.
Following EVO (Nguyen et al., 2024a) and protein language models (Lin et al., 2023b), we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness (Notin et al., 2022). EVO obtained (DMS) datasets with bacterial (prokaryote) and human (eukaryote) proteins from ProteinGYM at https://proteingym.org. To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies (Notin et al., 2022; Altae-Tran et al., 2021). For generative pre-trained models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations. The model performances of zero-shot function prediction are measured by the strength of Spearman’s Rank Correlation Coefficient (SRCC) that correlates the predicted likelihoods with the experimental fitness measurements. Full results of protein fitness prediction are shown in Table 5 and Table A6, in which our Life-Code achieves balancing performances with bacterial and human proteins and surpasses NT-2500M and EVO-7B.
ncRNA-Protein Interactions.
Following LucaOne (He et al., 2024), we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins. Since experimentally identifying ncRPI) is typically expensive and time-consuming, the AI-based ncRPI can be a promising task. LucaOne proposes a binary classification task involving pairs of ncRNA and Amino Acid sequences (20,824 pairs in total) with top-1 accuracy as the metric.
Central Dogma Evaluation.
To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne (He et al., 2024) to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated. LucaOne collects a total of 8,533 accurate DNA-protein pairs from 13 species in the NCBI RefSeq database, with each DNA sequence extending to include an additional 100 nucleotides at both the 5’ and 3’ contexts, preserving intron sequences within the data. LucaOne generated double the number of negative samples by implementing substitutions, insertions, and deletions within the DNA sequences or altering amino acids in the protein sequences to ensure the resultant DNA sequences could not be accurately translated into their respective proteins.

### v2

@@
 The interactions between DNA, RNA, and proteins are fundamental to biological processes, as illustrated by the central dogma of molecular biology.
-While modern biological pre-trained models have achieved great success in analyzing these macromolecules individually, their interconnected nature remains under-explored.
-In this paper, we follow the guidance of the central dogma to redesign both the data and model pipeline and offer a comprehensive framework, Life-Code, that spans different biological functions.
+Although modern biological pre-trained models have achieved great success in analyzing these macromolecules individually, their interconnected nature remains underexplored.
+This paper follows the guidance of the central dogma to redesign both the data and model pipeline and offers a comprehensive framework, Life-Code, that spans different biological functions.
 As for data flow, we propose a unified pipeline to integrate multi-omics data by reverse-transcribing RNA and reverse-translating amino acids into nucleotide-based sequences.
-As for the model, we design a codon tokenizer and a hybrid long-sequence architecture to encode the interactions of both coding and non-coding regions with masked modeling pre-training.
+As for the model, we design a codon tokenizer and a hybrid long-sequence architecture to encode the interactions between coding and non-coding regions through masked modeling pre-training.
 To model the translation and folding process with coding sequences, Life-Code learns protein structures of the corresponding amino acids by knowledge distillation from off-the-shelf protein language models.
 Such designs enable Life-Code to capture complex interactions within genetic sequences, providing a more comprehensive understanding of multi-omics with the central dogma.
-Extensive Experiments show that Life-Code achieves state-of-the-art performance on various tasks across three omics, highlighting its potential for advancing multi-omics analysis and interpretation.
-Machine Learning, ICML
+Extensive experiments show that Life-Code achieves state-of-the-art results on various tasks across three omics, highlighting its potential for advancing multi-omics analysis and interpretation.
 1 Introduction
-The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins (Smith & Doe, 2023; Li & Wang, 2024; Li et al., 2024; Liu et al., 2024a).
+Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM) severally proposed in each omics.
+The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins smith2023largescale ; liwang2024protein ; icml2024vqdna ; liu2024genbench .
 These endeavors aim to leverage deep learning to uncover functional elements, predict gene regulation, and accelerate protein engineering.
-Notably, the successful application of Transformer-based architectures in natural language processing (NLP) (Vaswani et al., 2017; Devlin et al., 2019) has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction (Zhou et al., 2022), protein structure elucidation (Jumper et al., 2021a), and more.
-Table 1: Overview of Bio-Foundation Models.
-Unlike DNA models NT-2500M or EVO-7B using BERT or generative pre-training with pure DNA sequences, our Life-Code models the Central Dogma with the unified nucleotide input format.
-We consider four aspects of evaluations: The DNA tasks report the averaged MCC (%) over the Promoter Detection (all) and Covid Variants Classification, the RNA task considers the Splicing Site Reconstructed in the GUE benchmark, the Protein tasks report the averaged SRCC (%) Protein Fitness Prediction with DMS, and the Central Dogma evaluation considers the translation between CDS and Amino Acids with top-1 accuracy (%).
-Full results in Table A6.
-Method NT-2500M EVO-7B LucaOne Life-Code
-# Params 2500M 7B 110M 350M
-Learning Tasks MLM MLM MLM+SL MLM+KD
-Data Source DNA DNA Multi-omics DNA+AA
-DNA Task 82.0 73.6 83.3 82.6
-RNA Task 89.4 87.5 89.1 89.8
-Protein Task 7.1 28.2 32.9 34.1
-Central Dogma -- 75.5 84.8 85.6
-Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM) severally proposed in each omics.
+Notably, the successful application of Transformer-based architectures in natural language processing (NLP) vaswani2017attention ; devlin2019bert has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction zhou2022transformer , protein structure elucidation jumper2021alphafold , and more.
+Despite the promise of these paradigms of foundation models in computational biology lin2023esm ; nguyen2024evo ; shen2024rnafm , three major obstacles continue to hamper multi-omics modeling: (1) Data Island.
+Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins hong2022crossomics , as shown in Figure 1.
+As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized.
+(2) Naïve “Copy and Paste”.
+While Transformers vaswani2017attention ; devlin2019bert have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability rives2021biological .
+For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes.
+(3) Inefficiency for Long Sequences.
+Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length alquraishi2019proteinnet .
+This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
 Figure 2: Illustration of Life-Code Framework, which contains three pipelines.
@@
 (c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning (SFT).
-Despite the promise of these paradigms of foundation models in computational biology (Lin et al., 2023b; Nguyen et al., 2024a; Shen et al., 2024), three major obstacles continue to hamper multi-omics modeling: (1) Data Island.
-Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins (Hong et al., 2022), as shown in Figure 1.
-As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized.
-(2) Naïve “Copy and Paste”.
-While Transformers (Vaswani et al., 2017; Devlin et al., 2019) have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability (Rives et al., 2021).
-For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes.
-(3) Inefficiency for Long Sequences.
-Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length (AlQuraishi, 2019).
-This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
 This paper introduces Life-Code, as shown in Figure 2, a unifying framework that addresses these limitations through multi-omics unification, bio-interpretable designs, and long-sequence efficiency.
@@
 This hybrid strategy yields significant computational savings over naïve $\mathcal{O}(n^{2})$ Transformer approaches.
-  • Comprehensive Validations.
-We evaluate Life-Code on diverse tasks in DNA, RNA, and protein applications.
-Empirical results confirm the effectiveness of our unified approach in these three fundamental modalities.
+  • Comprehensive Validations. Evaluated on diverse tasks in DNA, RNA, and protein applications, empirical results confirm the generalizability of Life-Code in these fundamental modalities.
 2 Method
@@
 Linear or sparse approximations (e.g., Mamba) reduce complexity but often sacrifice representational fidelity for complex biological contexts.
-Hence, the recent efficient hybrid strategy (Yang et al., 2025, 2024; Liu et al., 2024b, c) should be considered.
+Hence, the recent efficient hybrid strategy yang2025deltanet ; yang2024gateddelta ; iclr2024MogaNet ; icml2024chela ; ijcai2024longvq should be considered.
 We now formally describe how we unify DNA, RNA, and protein sources into nucleotides.
-Figure 3: Data Sampling Pipeline.
-During pre-training, we sample the DNA sequences and the pair of CDS and amino acids from the DNA dataset and the DNA-AA paring dataset.
-Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample.
-As for the DNA sequence, we employ truncation and padding.
-As for the CDS and its corresponding amino acids, we apply the packing strategy (Warner et al., 2024) to sample with several CDSs.
 2.2 Data Pipeline for Life-Code
@@
 We choose the canonical codon (or the most frequent codon) for each amino acid, yielding a surjective function $g:\Sigma\to\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}^{3}.$ Thus, a protein of length $L$ becomes a DNA-like sequence of length $3L$.
+Figure 3: Data Sampling Pipeline.
+During pre-training, we sample the DNA sequences and the pair of CDS and amino acids from the DNA dataset and the DNA-AA pairing dataset.
+Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample.
+As for the CDS and its corresponding amino acids, we apply the packing strategy Warner2024ModernBERT to sample with several CDSs.
 Data sampling.
-Following GenBank (Benson et al., 2012) curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones.
+Following GenBank benson2012genbank curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones.
 For CDS–Amino Acid pairs, we extract annotated CDS records and their corresponding protein sequences, then pack multiple short segments into a single training sample with special separators.
@@
 To model the interactions among DNA, RNA, and Amino Acids, our tokenizer takes nucleotide acids (4 words) as the inputs, then takes codons as the latent vocabulary (64 words), and translates them to Amino Acids (20 words) as the outputs, which could be pre-trained by (a) nucleotide masked modeling and (b) CDS to Amino Acid translation.
-Figure 5: Illustration of Life-Code Encoder pre-training pipeline.
-We model the long sequence by the hybrid encoder as a mixture of RNN blocks (DeltaNet) and self-attention blocks.
-With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma (with span masks), and (c) the knowledge distillation (KD) with pre-trained protein LM.
-Note that red lines indicates the learning objectives.
 2.3 Life-Code Tokenizer Training
@@
 Bi-directional input.
-To use the double-helix property as Schiff et al. (2024a), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
+To use the double-helix property as Caduceus (icml2024caduceus, ), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
 $(\mathbf{E}^{(+)},\ \mathbf{E}^{(-)})=\mathrm{Split}(\mathbf{E}),$
@@
 […]
-yielding a representation that encodes both the forward and reverse complement features, where $\mathbf{H}\in\mathbb{R}^{D}$.
+where $\mathbf{H}\in\mathbb{R}^{D}$ encodes both the forward and reverse complement features.
 Hybrid Token Mixer.
-Inspired by recent advances in EVO-style architectures (Nguyen et al., 2024a) and Gated Delta Networks (GDN) (Yang et al., 2024), we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA).
+Inspired by recent advances in EVO-style architectures nguyen2024evo and Gated Delta Networks (GDN) yang2024gateddelta , we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA).
 Concretely, for every twelve encoder layers, eleven employ the GDN update rule—offering linear or near-linear complexity via kernel-based reordering—while one layer applies MHSA.
@@
 This 11 : 1 ratio of GDN to dense layers significantly boosts computational efficiency (compared to a pure $\mathcal{O}(n^{2})$ Transformer) while preserving model expressiveness and long-distance modeling capacity.
+Encoder Pre-training.
+Once the decoder and translator of our tokenizer weights are fixed (see Section 2.3), we train the Life-Code Encoder (Figure 5) on three tasks: (1) Masked DNA Reconstruction (MLM), (2) CDS-to-Amino-Acid Translation, and (3) Token-Level Knowledge Distillation (KD) from a pretrained Protein LM.
+The first two objectives match those used to pre-train the tokenizer but are now applied to the encoder output.
+Specifically, the encoder generates codon embeddings that a DNA De-tokenizer can reconstruct (MLM) and that an Amino Acid Translator can convert to protein residues (Translation).
+For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues).
+Inspired by the DINO framework (oquab2024dinov2, ), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (lin2022ESM2, ).
+Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 4D\rightarrow D^{\prime}$.
+Then, the alignment or KD loss is defined as:
+$\mathcal{L}_{\mathrm{KD}}\;=\;\sum_{j=1}^{L}-P_{\theta^{\prime}}\bigl{(}\mathbf{h}_{j}^{\mathrm{PLM}}\bigr{)}^{\top}\log P_{\theta}\bigl{(}\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}})\bigr{)}.$ (4)
+We combine these three losses using a log-sum multi-task strategy (lin2023dualmtl, ) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
+$\log\bigl{(}\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr{)},$ (5)
+where $\epsilon$ is a small constant.
+This adaptive weighting prevents any single task from dominating and ensures stable optimization across various training stages.
+Notably, we also adopt this strategy when training the tokenizer.
+Figure 5: Illustration of Life-Code Encoder pre-training pipeline.
+We model the long sequence by the hybrid encoder as a mixture of RNN blocks (DeltaNet) and self-attention blocks.
+With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma (with span masks), and (c) the knowledge distillation (KD) with pre-trained protein LM.
+Note that red lines indicates the learning objectives.
 Figure 6: Empirical Analysis of Life-Code Tokenizer.
@@
 Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons (black crosses) form a distinct region.
-These views jointly demonstrate that Life-Code encodes both species-dependent codon usage preferences and biologically meaningful codon–amino acid relationships.
-Encoder Pre-training.
-Once the decoder and translator of our tokenizer weights are fixed (see Section 2.3), we train the Life-Code Encoder (Figure 5) on three tasks: (1) Masked DNA Reconstruction (MLM), (2) CDS-to-Amino-Acid Translation, and (3) Token-Level Knowledge Distillation (KD) from a pretrained Protein LM.
-The first two objectives match those used to pre-train the tokenizer but are now applied to the encoder output.
-Specifically, the encoder generates codon embeddings that a DNA De-tokenizer can reconstruct (MLM) and that an Amino Acid Translator can convert to protein residues (Translation).
-For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues).
-Inspired by the DINO framework (Oquab et al., 2024), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (Lin et al., 2022).
-Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 2D\rightarrow D^{\prime}$.
-Then, the alignment or KD loss is defined as:
-$\mathcal{L}_{\mathrm{KD}}\;=\;\sum_{j=1}^{L}-P_{\theta^{\prime}}\bigl{(}\mathbf{h}_{j}^{\mathrm{PLM}}\bigr{)}^{\top}\log P_{\theta}\bigl{(}\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}})\bigr{)}.$ (4)
-We combine these three losses using a log-sum multi-task strategy (Lin et al., 2023a) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
-$\log\bigl{(}\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr{)},$ (5)
-where $\epsilon$ is a small constant.
-This adaptive weighting prevents any single task from dominating and ensures stable optimization across various training stages.
-Notably, we also adopt this strategy when training the tokenizer.
+These views jointly demonstrate that Life-Code encodes both species-dependent codon usage preferences and biologically meaningful relationships.
 2.5 Empirical Consequences
@@
 This structural arrangement confirms that the model internalizes codon–amino acid mappings, reflecting both frequency patterns and higher-level attributes like polarity or aromaticity.
-Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals—ranging from species-specific usage preferences to fundamental amino acid groupings.
+Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals—ranging from species-specific usage preferences.
 Figure 7: Training Efficiency on a Single A100 GPU.
 We measure throughput (K t/s) at four sequence lengths–batch size settings.
-DNABERT-2 (blue) slows significantly at higher lengths and reaches an out-of-memory (OOM) error at $16\text{K}\times 2$.
-Both EVO (red) and Life-Code (green) scale more effectively, with Life-Code consistently achieving the highest throughput.
-Table 2: Comparison on Genomic Benchmarks.
-Top-1 accuracy (%) averaged over several similar tasks is reported for popular DNA foundation models with SFT evaluation, where the best and the second best results are marked as the bold and underlined types.
-Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
-# Params (M) 6.6 86 117 113 498 7.9 93 100 1.7 350
-Model Type RNN BERT BERT BERT BERT RNN BERT BERT CNN Hybrid
-Enhancers (3 tasks) 80.88 80.14 82.81 83.22 84.56 79.96 82.37 82.79 80.90 85.27
-Species Classification (2 tasks) 93.61 94.74 95.49 95.11 96.64 94.65 95.79 96.46 95.50 96.95
-Regulatory Elements (3 tasks) 88.89 83.42 86.33 87.89 89.05 85.97 87.62 90.57 87.30 90.87
-Table 3: Comparison on GUE Benchmark.
-MCC (%) averaged across several sub-tasks is reported with SFT evaluation for seven primary tasks, including Epigenetic Marks Prediction, Human Transcription Factor (TF) Prediction, Mouse Transcription Factor Prediction, Core Promoter Detection, Promoter Detection, Splice Site Reconstructed, and Covid Variants Classification (Virus Covid).
-The best and the second best results are marked as the bold and underlined types.
-Method HyenaDNA DNABERT NT-2500M-multi DNABERT2 Caduceus-PS VQDNA MxDNA ConvNova Life-Code
-# Params (M) 6.6 86 2537 117 1.9 93 100 1.7 350
-Epigenetic Marks Prediction (10 tasks) 58.94 49.08 58.06 55.98 58.39 57.95 67.29 68.91 65.95
-Human TF Detection (3 tasks) -- 64.17 63.34 70.11 -- 70.56 -- -- 71.26
-Mouse TF Detection (3 tasks) -- 56.43 67.02 67.99 -- 69.80 -- -- 71.56
-Core Promoter Detection (3 tasks) -- 71.81 71.63 70.53 -- 73.37 -- -- 73.51
-Promoter Detection (3 tasks) -- 81.69 88.15 84.21 -- 86.58 -- -- 88.44
-Splice Site Reconstructed -- 84.07 89.35 84.99 -- 89.53 -- -- 89.76
-Virus Covid Classification -- 55.50 73.04 71.02 -- 74.32 -- -- 73.82
+DNABERT-2 slows significantly at higher lengths and reaches an out-of-memory (OOM) at $16\text{K}\times 2$.
 Training efficiency
@@
 Similarly, we also design the DNA De-Tokenizer and Amino Acid Translator for tokenizer pre-training.
-As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (Loshchilov & Hutter, 2019) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2.
-As shown in Table A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases as described in Figure A1 and Appendix A.
-View Appendix A.2 for details.
+As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (iclr2019AdamW, ) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2.
+As shown in Table A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases, as described in Figure A1, Appendix A, and  A.2.
 Life-Code Encoder.
-Similar to the BERT-Large architecture (Devlin et al., 2019), the Life-Code encoder has 24 layers in total with an embedding dim of 1024 and 340M parameters.
-We enhance the model with three-fold designs: (i) Mixture of DeltaNet (Yang et al., 2025) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block.
-(ii) The LLaMA-like macro design (Touvron et al., 2023) with RMSNorm (Zhang & Sennrich, 2019), Layer Scale (Touvron et al., 2021), Rotary Position Embedding (RoPE) (Su et al., 2021), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences.
-(iii) During pre-training, we apply the packing strategy (Warner et al., 2024) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences.
+Similar to the BERT-Large architecture (devlin2019bert, ), the Life-Code encoder has 24 layers in total with an embedding dim of 1024/1280 and 340/500M parameters.
+We enhance the model with three-fold designs: (i) Mixture of DeltaNet (yang2025deltanet, ) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block.
+(ii) The LLaMA-like macro design (Touvron2023LLaMA, ) with RMSNorm (Zhang2019RMSNorm, ), Layer Scale (iccv2021CaiT, ), Rotary Position Embedding (RoPE) (Su2021RoFormer, ), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences.
+(iii) During pre-training, we apply the packing strategy (Warner2024ModernBERT, ) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences.
 With three pre-training tasks in Eq.
-5, the Life-Code tokenizer and encoder are optimized by AdamW optimizer for 1M iterations with the batch size of 256 and the basic learning rate of $1\times 10^{-4}$.
+5, the Life-Code tokenizer and encoder are optimized by the AdamW optimizer for 1M iterations with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$.
 We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation.
-As for KD from a pre-trained Protein LM, we adopted ESM2-650M (Lin et al., 2022) and a protein decoder with an output dimension of 1280.
+As for KD from a pre-trained Protein LM, we adopted ESM2-650M (lin2022ESM2, ) and a protein decoder with an output dimension of 1280.
 During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations.
@@
 View Appendix A.2 for more details.
+Table 1: Comparison on Genomic Benchmarks.
+Top-1 accuracy (%) averaged over several similar tasks is reported for popular DNA foundation models with SFT evaluation.
+The best and the second best results are marked as the bold and underlined types.
+Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
+# Params (M) 6.6 86 117 113 498 7.9 93 100 1.7 500
+Model Type RNN BERT BERT BERT BERT RNN BERT BERT CNN Hybrid
+Enhancers (3 tasks) 80.88 80.14 82.81 83.22 84.56 79.96 82.37 82.79 80.90 86.00
+Species Classification (2 tasks) 93.61 94.74 95.49 95.11 96.64 94.65 95.79 96.46 95.50 97.20
+Regulatory Elements (3 tasks) 88.89 83.42 86.33 87.89 89.05 85.97 87.62 90.57 87.30 91.30
+Average 87.07 85.02 87.30 87.94 89.26 85.89 87.69 89.12 86.95 90.79
+Table 2: Comparison on GUE Benchmark.
+MCC (%) averaged across several sub-tasks is reported with SFT evaluation for seven primary tasks (24 sub-tasks in total).
+The best and the second best results are marked as the bold and underlined types.
+Method HyenaDNA DNABERT NT-multi DNABERT2 Caduceus-PS VQDNA MxDNA ConvNova Life-Code
+# Params (M) 6.6 86 2537 117 1.9 93 100 1.7 500
+Epigenetic Marks Prediction (10) 58.94 49.08 58.06 55.98 58.39 57.95 67.29 68.91 67.85
+Human TF Detection (3) -- 64.17 63.34 70.11 -- 70.56 -- -- 71.34
+Mouse TF Detection (3) -- 56.43 67.02 67.99 -- 69.80 -- -- 71.76
+Core Promoter Detection (3) -- 71.81 71.63 70.53 -- 73.37 -- -- 74.53
+Promoter Detection (3) -- 81.69 88.15 84.21 -- 86.58 -- -- 89.28
+Splice Site Reconstructed (1) -- 84.07 89.35 84.99 -- 89.53 -- -- 90.24
+Virus Covid Classification (1) -- 55.50 73.04 71.02 -- 74.32 -- -- 74.88
+Average (24 tasks) 58.94 60.53 67.23 66.43 58.39 68.51 67.29 68.91 73.51
 3.2 Comparison on Genomic Benchmarks
-As DNA sequences are the foundation of the central dogma, we primarily compare with latest DNA foundation models with the SFT evaluation protocol proposed by (Nguyen et al., 2024c; Zhou et al., 2024), following benchmark evaluations for fair comparison.
-We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (Ji et al., 2021a), DNABERT2 (Zhou et al., 2024), GENA-LM (Fishman et al., 2023), Nucleotide Transformer (NT) (Dalla-Torre et al., 2023), with different tokenizers like VQDNA (Li et al., 2024) and MxDNA (Qiao et al., 2024); (b) RNN-based linear attentions including HyenaDNA (Nguyen et al., 2024c) and Caduceus (Schiff et al., 2024a); (c) CNN model ConvNova (Anonymous, 2025).
+As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna, ; iclr2024dnabert2, ), following benchmark evaluations for fair comparison.
+We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert, ), DNABERT2 (iclr2024dnabert2, ), GENA-LM (fishman2023GenaLM, ), Nucleotide Transformer (NT) (NM2023NucleotideTrans, ), with different tokenizers like VQDNA (icml2024vqdna, ) and MxDNA (nips2024MXDNA, ); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna, ) and Caduceus (icml2024caduceus, ); (c) CNN model ConvNova (iclr2025convnova, ).
 Genomic Benchmarks.
-We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (Grešová et al., 2023), which we adopted results from GenBench implementations (Liu et al., 2024a) and all compared DNA models are fully SFT and reported the top-1 accuracy.
-Table 2 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages of both RNN and self-attention blocks.
+We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark, ), which we adopted results from GenBench implementations (liu2024genbench, ), and all compared DNA models are fully SFT and reported the top-1 accuracy.
+Table 1 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages.
 GUE Benchmarks.
-We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (Zhou et al., 2024), as shown in Table 3, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification.
-All results are adopted from DNABERT2 or each paper with their official SFT setups, and Matthews Correlation Coefficient (MCC) is reported as the metric.
-Compared to popular DNA models, our Life-Code shows superiority in gene annotation tasks, which are important regulation processes in the central dogma, while achieving competitive results in Epigenetic regulating tasks.
-Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator.
-Top-1 AUC score or F1 score are reported for DNA models or RNA models with SFT evaluation, respectively.
-The best and the second best results are marked as the bold and underlined types.
-Dataset SpliceAI DNABERT2 NT Caduceus Life-Code
-SpliceAI 63.2 67.1 63.9 69.1 69.6
-Dataset Spliceator DDSP RNA-FM RINALMo Life-Code
-Spliceator 91.0 91.9 92.1 94.8 94.5
+We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2, ), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification.
+All results are adopted from DNABERT2 or each paper with their official SFT setups, and the Matthews Correlation Coefficient (MCC) is reported as the metric.
 3.3 Comparison on RNA Splicing Site Prediction
@@
 We consider two typical datasets with the SFT evaluation protocol.
-As for SpliceAI dataset (Jaganathan et al., 2019), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma.
-As for Spliceator dataset (Scalzitti et al., 2021), we then consider RNA splicing methods (Spliceator, DDSP (Naito, 2019), RINALMo (Penic et al., 2024)) and RNA language model, RNA-FM (Shen et al., 2024), with RNA data of multi-species.
-As shown in Table 4, our Life-Code achieves competitive F1 scores with the specially designed splicing method and surpasses RNA-FM.
-Table 5: Comparison on Zero-shot Protein Fitness Prediction.
-SRCC (%) is reported on DMS datasets.
-The best and the second best results are marked as the bold and underlined types.
-ESM2 is marked as gray since it takes the amino acids as the input, while other models utilize the gene sequences.
-Method GenSLM NT-2500M EVO ESM2-650M Life-Code
-# Params 2.5B 2.5B 7B 650M 350M
-Bacteria 24.7 9.4 45.3 51.2 45.7
-Human 6.9 4.7 11.1 37.5 22.4
+As for the SpliceAI dataset (JAGANATHAN2019SpliceAI, ), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma.
+As for Spliceator dataset (BMC2021spliceator, ), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP, ), RINALMo (Penic2024RiNALMo, )) and RNA language model, RNA-FM (shen2024rnafm, ), with RNA data of multi-species.
+As shown in Table 4, our Life-Code achieves the best F1 scores with the specified baselines.
+Table 3: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets.
+Method GenSLM NT-2500M EVO Life-Code
+# Params 2.5B 2.5B 7B 500M
+Bacteria 24.7 9.4 45.3 46.4
+Human 6.9 4.7 11.1 23.8
+Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
+Dataset SpliceAI DNABERT2 NT Caduceus Life-Code
+SpliceAI 63.2 67.1 63.9 69.1 70.4
+Dataset Spliceator DDSP RNA-FM RINALMo Life-Code
+Spliceator 91.0 91.9 92.1 94.8 95.0
 3.4 Comparison on Protein Fitness Prediction
-Following EVO (Nguyen et al., 2024a), we then evaluate foundation models pre-trained with DNA or protein data with Protein Fitness Prediction tasks using zero-shot evaluation protocol (without task-specific SFT).
-Deep Mutational Scanning (DMS) datasets (Notin et al., 2022; Nguyen et al., 2024a) introduce a comprehensive set of mutations into protein CDS to measure their effects on fitness, where fitness serves as a metric for how effectively a protein performs a specific function.
-As shown in Table 5, bacterial (prokaryote) and human (eukaryote) protein fitness is evaluated by Spearman’s Rank Correlation Coefficient (SRCC), where the protein specialist model ESM2 achieves the best result.
-Life-Code outperforms other DNA models like GenSLM (Zvyagin et al., 2022) and EVO-7B with fewer parameters, which might be attributed to its modeling of the translation and folding processes in the central dogma.
-Table 6: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation. Bold type denotes the best result.
+Following EVO (nguyen2024evo, ), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction.
+Deep Mutational Scanning (DMS) datasets (icml2022Tranception, ; nguyen2024evo, ) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy.
+For both bacterial and human proteins, we compute Spearman’s rank correlation (SRCC) between model scores and experimental measurements (Table 3).
+The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms, ) and the 7-B-parameter EVO model—despite using fewer parameters.
+We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without any task-specific fine-tuning.
+Table 5: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation.
 Method ncRNA-Protein Central Dogma
@@
 LucaOne 93.80 84.84
-Life-Code 93.45 85.63
-3.5 Comparison on Multi-omics Tasks
-We conduct further comparisons with proposed biological foundation models (EVO (Nguyen et al., 2024a) and LucaOne (He et al., 2024)) with multi-omics tasks.
-Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins.
-As shown in Table 6, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve the competitive result as the supervised LucaOne, which should process two inputs with separate modules.
-Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and the protein sequences, where Life-Code achieves the best accuracy over existing models.
+Life-Code 94.12 86.51
 Table 6: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory on the single A100-80G.
@@
 DNABERT2 117M 61.0 86.8 99.3 OOM OOM
-Life-Code 350M 62.2 87.6 99.3 99.4 99.5
+Life-Code 500M 62.8 88.1 99.3 99.4 99.5
+Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss; 3:1 exceeds memory limits.
+3.5 Comparison on Multi-omics Tasks
+We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo, ) and LucaOne (he2024lucaone, )) with multi-omics tasks.
+Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins.
+As shown in Table 5, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve a competitive result as the supervised LucaOne, which should process two inputs with separate modules.
+Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy over existing models.
 3.6 Empirical Analysis
-After verifying the effectiveness with extensive comparison results, we further consider long-sequence tasks like HyenaDNA (Nguyen et al., 2024c) in addition to Figure 7.
-We compare the inference throughput of our bi-directional hybrid model to baselines (FlashAttention and HyenaDNA).
-In Table 7, empirical profiling shows that LifeCode achieves competitive speed with these specialized methods while retaining strong global modeling capacity.
-In particular, mixing a small fraction of MHSA enables better downstream performance without incurring a large computational penalty, e.g., splicing site prediction in Table 4.
+Figure 8 shows that a pure RNN stack (12:0) is fastest ( $\sim$45 K token/s) but sacrifices 2–3 pp of accuracy.
+Injecting just one MHSA layer (11:1) restores nearly all lost performance—DNA 95 to 97 %, Protein 32 to 34 %, CD 83 to 86 %—while reducing throughput by only $\sim$9 %.
+A 5:1 mix boosts accuracy further with a modest extra cost, whereas 3:1 exhausts memory.
+Thus, adding a small fraction of MHSA markedly improves downstream accuracy without a large computational penalty.
 Table 7: Ablation Study of Life-Code tokenizer, encoder, and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
@@
 DNA Foundation Models.
-Recent years have seen rapid progress in DNA foundation models that leverage Transformer-like architectures.
-Early attempts such as DNABERT (Ji et al., 2021b) treated genomic sequences akin to linguistic tokens, enabling pre-trained contextual representations.
-Follow-up work DNABERT-2 (Zhou et al., 2023) expanded this paradigm by introducing more efficient training protocols and supporting multi-species data.
-In parallel, The Nucleotide Transformer (Dalla-Torre et al., 2023) demonstrated the feasibility of scaling up Transformer architectures for human genomics.
-Beyond Transformers, kernel-based or hierarchical approaches emerged—e.g., HyenaDNA (Nguyen et al., 2024b), which reduces the quadratic complexity of attention for extremely long sequences.
-Meanwhile, alternative designs like Caduceus (Schiff et al., 2024b) incorporate selective structural priors for long-range DNA modeling, and MxDNA (Qiao et al., 2024) explores adaptive tokenization schemes that automatically discover suitable patterns for genomic data.
+Recent DNA foundation models have evolved swiftly from DNABERT ji2021dnabert , which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 zhou2023dnabert2 and the large-scale Nucleotide Transformer NM2023NucleotideTrans .
+To tame quadratic attention on longer sequences, HyenaDNA nguyen2024hyenadna replaces full self-attention with efficient hierarchical kernels.
+Other variants embed biological priors—Caduceus schiff2024caduceus for long-range structure and MxDNA nips2024MXDNA for adaptive tokenization—illustrating a broad shift toward specialized, scalable architectures for genomics.
 Protein and RNA Advances.
-On the protein side, AlphaFold series (Jumper et al., 2021a, b) revolutionized structure prediction, catalyzing a surge in protein-based language models such as ESM-2 (Lin et al., 2022), which refines large-scale protein embeddings.
-Beyond proteins, RNA modeling also gained traction; for instance, (Shen et al., 2024) employed a language-model-based technique for accurate 3D structure prediction.
-These efforts underscore the trend toward specialized architectures for each biomolecule yet also highlight the desire for integrative multi-omics solutions.
+AlphaFold jumper2021alphafold ; jumper2021alphafold2 ignited the protein-LM wave, inspiring models like ESM-2 lin2022ESM2 .
+RNA efforts followed, e.g., RNA-FM shen2024rnafm for 3D structure prediction.
+While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals prakash2024bridging ; bioinfo2024genomic ; BioInfo2023NTA ; nmil2024CaLM , delivering notable downstream gains Ren2024COMET and pointing toward unified multi-omics models.
 Multi-Omics Modeling.
 Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks.
-CD-GPT (Zhu et al., 2024) explicitly connects DNA, RNA, and proteins through the central dogma, while BSM (Xiang et al., 2024) highlights the potential for small but effective models covering genes and proteins simultaneously.
-Evo (Nguyen et al., 2024a) similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach.
-Closely related is LucaOne (He et al., 2024), which advocates a unified nucleic acid and protein language for biological representation.
+CD-GPT zhu2024CDGPT explicitly connects DNA, RNA, and proteins through the central dogma, while BSM xiang2024BSM highlights the potential for small but effective models covering genes and proteins simultaneously.
+EVO nguyen2024evo similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach.
+Closely related is LucaOne he2024lucaone , which advocates a unified nucleic acid and protein language for biological representation.
 5 Conclusion and Limitations
-We have introduced Life-Code, a method that unifies DNA, RNA, and protein sequences by mapping the latter two back to a DNA-like representation.
-This design offers straightforward multi-omics integration and maintains biological interpretability via differential encoding of coding versus non-coding regions.
-To handle the long-range dependencies inherent in genomic data, we employ an efficient symmetric convolution-based architecture.
-Moreover, knowledge distillation from large protein models allows our approach to scale without excessive resource demands.
+Life-Code unifies DNA, RNA, and protein by remapping all sequences to nucleotides, enabling seamless multi-omics integration while preserving biological meaning (coding vs non-coding).
+A symmetric convolution backbone captures long-range genomic context, and distilling knowledge from large protein LMs keeps training efficient.
 Limitations
@@
 Despite these challenges, we believe Life-Code offers a promising step toward comprehensive multi-omics data.
-Impact Statements
-This paper presents a novel approach, Life-Code, aimed at advancing the field of machine learning and bioinformatics by unifying the modeling of multi-omics data through the guidance of the Central Dogma of molecular biology.
-By integrating DNA, RNA, and protein data into a single framework, our work has the potential to improve the understanding of complex biological processes and drive advancements in genomics, proteomics, and systems biology.
-Ethical Aspects
-This work has significant potential applications in fields such as personalized medicine, drug discovery, and disease diagnostics.
-However, we acknowledge the need to address concerns related to data privacy and the responsible use of sensitive genetic information.
-We have adhered to ethical guidelines in our experiments, ensuring that no private or identifiable data was used.
 Appendix
@@
 Figure A1: Data Collection Pipeline.
-We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA paring datasets.
-Appendix A Implementation Details
+We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA pairing datasets.
 Table A1: Configuration of pre-training datasets.
@@
 mRNA-AA pair 3.908,074 GenBank
+Appendix A Implementation Details
 A.1 Pre-training Dataset
 We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1.
-As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer (Dalla-Torre et al., 2023) and DNABERT2 (Zhou et al., 2024).
+As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer [8] and DNABERT2 [64].
 As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma.
-We also collect some Amino Acids in the UniRef50 database following LucaOne (He et al., 2024) and obtain their corresponding cDNA by reverse translation with online tools.
+We also collect some Amino Acids in the UniRef50 database following LucaOne [13] and obtain their corresponding cDNA by reverse translation with online tools.
 We provide detailed information for used datasets in Table A1.
@@
 It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS).
-The pre-trained protein language model employs the amino acid vocabulary of ESM-2 (Lin et al., 2022).
+The pre-trained protein language model employs the amino acid vocabulary of ESM-2 [26].
 Table A2: Configuration of the network designs and pre-training settings for Life-Code models.
@@
 Configuration Tokenizer Encoder
-Embedding dim 384 1024
+Embedding dim 384 1024/1280
 Block number 1 24
@@
 Attention heads 0 24
-Parameters 8M 340M
+Parameters 8M 350/500M
 Optimizer AdamW
@@
 Weight decay $1\times 10^{-2}$
-Base learning rate $2\times 10^{-4}$ $1\times 10^{-4}$
-Batch size 512 256
+Base learning rate $4\times 10^{-4}$ $1\times 10^{-4}$
+Batch size 1024 256
 LR scheduler Cosine Annealing
@@
 Pre-training Settings.
-As shown in Figure 4, we pre-train the Life-Code tokenizer with the DNA de-tokenizer and Amino Acid de-tokenizer by AdamW optimizer for 100,000 iterations (randomly sampled datasets) with a basic learning rate of 2e-4 and a batch size of 512, as detailed in Table A2.
-We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 8 and a gradient accumulation time of 4.
+As shown in Figure 4, we pre-train the Life-Code tokenizer with the DNA de-tokenizer and Amino Acid de-tokenizer by AdamW optimizer for 100,000 iterations (randomly sampled datasets) with a basic learning rate of $2\times 10^{-4}$ and a batch size of 1024, as detailed in Table A2.
+We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 8 and a gradient accumulation time of 16.
 A.3 Life-Code Encoder
 Encoder Architecture.
-As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet (Yang et al., 2025)) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 (MiniMax et al., 2025), which could utilize the complementary properties of GDN and MHSA while maintaining efficiency.
-(2) The model macro design employs pre-norm (Wang et al., 2019a) with RMSNorm (Zhang & Sennrich, 2019), Layer Scale (Touvron et al., 2021), Rotary Position Embedding (RoPE) (Su et al., 2021), SwiGLU (Touvron et al., 2023), and FlashAttention implementations to facilitate training large-scale models stably with long sequences.
-(3) During pre-training, we apply the packing strategy (Warner et al., 2024) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
+As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet [60]) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 [32], which could utilize the complementary properties of GDN and MHSA while maintaining efficiency.
+Notice that the DeltaNet variants [59] introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block.
+(2) The model macro design employs pre-norm [55] with RMSNorm [61], Layer Scale [52], Rotary Position Embedding (RoPE) [51], SwiGLU [53], and FlashAttention implementations to facilitate training large-scale models stably with long sequences.
+(3) During pre-training, we apply the packing strategy [57] to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
 Pre-training Settings.
 As shown in Figure 5, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq.
-5 for 1M steps with the batch size of 256 and the basic learning rate of $1\times 10^{-4}$.
+5 for 1M steps with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$.
 We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation.
-As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D(Lin et al., 2022) and a protein decoder with the output dimension of 1280.
+As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D[26] and a protein decoder with the output dimension of 1280.
 During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations.
@@
 In most cases, we apply Supervised Fine-tuning (SFT) to transfer pre-trained models to downstream tasks.
-Following (Nguyen et al., 2024c; Zhou et al., 2024), adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy (Hu et al., 2021) is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32.
-For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate {1e-5, 5e-5, 1e-4}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench (Liu et al., 2024a).
+Following [36, 64], adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy [15] is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32.
+For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate $1\times 10^{-5}$, $5\times 10^{-5}$, $1\times 10^{-4}$}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench [28].
 Note that the maximum input length will be determined for different tasks since the sequence lengths of downstream tasks vary widely.
@@
 B.1 DNA Tasks with Genomics Benchmark
-As proposed by (Grešová et al., 2023), three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark.
+As proposed by [12], three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark.
 As for the enhancer prediction, three datasets are provided for identifying enhancer regions in the mouse or human genome.
@@
 As for the regulatory elements classification, three datasets are used for classifying sequences as regulatory regions based on Ensembl annotations, identifying open chromatin regions, or identifying non-TATA promoter regions in the human genome.
-We utilize the fully reproduced results of various DNA models in GenBench (Liu et al., 2024a).
+We utilize the fully reproduced results of various DNA models in GenBench [28].
 B.2 DNA Tasks with GUE Benchmark
-As proposed by DNABERT2 (Zhou et al., 2024), the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric.
+As proposed by DNABERT2 [64], the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric.
 To comprehensively evaluate the genome foundation models in modeling variable-length sequences, tasks with input lengths ranging from 70 to 1000 are selected.
@@
 Accurate detection aids in understanding gene regulation and disease mechanisms.
-The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) (Dreos et al., 2013).
+The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) [10].
 Meanwhile, we construct the non-promoter class with equal-sized randomly selected sequences outside of promoter regions but with TATA motif (TATA non-promoters) or randomly substituted sequences (non-TATA, non-promoters).
@@
 This task predicts human transcription factor binding sites (TFBS), crucial for gene expression regulation.
-Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) (Consortium et al., 2012b) are collected via the UCSC genome browser.
+Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) [6] are collected via the UCSC genome browser.
 TFBS sequences are 101-bp regions around peaks, while non-TFBS sequences match in length and GC content.
@@
 This prediction is crucial to understanding protein diversity and the implications of aberrant splicing in genetic disorders.
-The dataset (Scalzitti et al., 2021) consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome.
-As suggested by Ji et al.
-(2021a), existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality.
+The dataset [46] consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome.
+As suggested by [17], existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality.
 As such, we reconstruct the dataset by iteratively adding adversarial examples (unseen false positive predictions in the hold-out set) in order to make this task more challenging.
 Transcription Factor Binding Site Prediction (Mouse).
-This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) (Consortium et al., 2012a) from the UCSC genome browser.
+This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) [7] from the UCSC genome browser.
 Negative examples are created by di-nucleotide shuffling.
@@
 This task aims to predict the variant type of the SARS_CoV_2 virus based on 1000-length genome sequences.
-We download the genomes from the EpiCoV database (Khare et al., 2021) of the Global Initiative on Sharing Avian Influenza Data (GISAID).
+We download the genomes from the EpiCoV database [21] of the Global Initiative on Sharing Avian Influenza Data (GISAID).
 We consider 9 types of SARS_CoV_2 variants, including Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda and Zeta.
@@
 B.3 mRNA Splicing Tasks
-Following (Zhou et al., 2024; Shen et al., 2024), we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression.
+Following [64, 49], we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression.
 During splicing, introns are removed from precursor messenger RNAs (pre-mRNAs), and exons are joined together to form mature mRNAs.
@@
 Identifying splice sites—the donor sites at the 5’ end of introns and the acceptor sites at the 3’ end—is vital for accurately predicting gene structure and location.
-Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset (Wang et al., 2019b) used in the GUE benchmark.
+Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset [56] used in the GUE benchmark.
 Spliceator dataset.
-This dataset (Scalzitti et al., 2021) consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans.
+This dataset [46] consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans.
 The gold standard dataset GS 1 is adopted, which contains an equal number of positive and negative samples, and the F1 score is used as the evaluation metric.
@@
 SpliceAI dataset.
-This dataset (Jaganathan et al., 2019) also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information.
+This dataset [16] also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information.
 SpliceAI also references the ClinVar database to evaluate the clinical significance of predicted splicing variants, which contains information on clinically relevant variants and their associations with diseases.
@@
 Zero-shot Protein Fitness Prediction.
-Following EVO (Nguyen et al., 2024a) and protein language models (Lin et al., 2023b), we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness (Notin et al., 2022).
+Following EVO [35] and protein language models [27], we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness [38].
 EVO obtained (DMS) datasets with bacterial (prokaryote) and human (eukaryote) proteins from ProteinGYM at https://proteingym.org.
-To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies (Notin et al., 2022; Altae-Tran et al., 2021).
+To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies [38, 2].
 For generative pre-trained models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations.
@@
 ncRNA-Protein Interactions.
-Following LucaOne (He et al., 2024), we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins.
+Following LucaOne [13], we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins.
 Since experimentally identifying ncRPI) is typically expensive and time-consuming, the AI-based ncRPI can be a promising task.
@@
 Central Dogma Evaluation.
-To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne (He et al., 2024) to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated.
+To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne [13] to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated.
 LucaOne collects a total of 8,533 accurate DNA-protein pairs from 13 species in the NCBI RefSeq database, with each DNA sequence extending to include an additional 100 nucleotides at both the 5’ and 3’ contexts, preserving intron sequences within the data.

### v3

@@
 1 Introduction
-Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM) severally proposed in each omics.
-The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins smith2023largescale ; liwang2024protein ; icml2024vqdna ; liu2024genbench .
+The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins (liwang2024protein; icml2024vqdna; liu2024genbench).
 These endeavors aim to leverage deep learning to uncover functional elements, predict gene regulation, and accelerate protein engineering.
-Notably, the successful application of Transformer-based architectures in natural language processing (NLP) vaswani2017attention ; devlin2019bert has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction zhou2022transformer , protein structure elucidation jumper2021alphafold , and more.
-Despite the promise of these paradigms of foundation models in computational biology lin2023esm ; nguyen2024evo ; shen2024rnafm , three major obstacles continue to hamper multi-omics modeling: (1) Data Island.
-Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins hong2022crossomics , as shown in Figure 1.
+Notably, the successful application of Transformer-based architectures in natural language processing (NLP) (vaswani2017attention; devlin2019bert) has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction (zhou2022transformer), protein structure elucidation (jumper2021alphafold), and more.
+Despite the promise of these paradigms of foundation models in computational biology (lin2023esm; nguyen2024evo; shen2024rnafm), three major obstacles continue to hamper multi-omics modeling: (1) Data Island.
+Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins (hong2022crossomics), as shown in Figure 1.
 As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized.
 (2) Naïve “Copy and Paste”.
-While Transformers vaswani2017attention ; devlin2019bert have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability rives2021biological .
+While Transformers (vaswani2017attention; devlin2019bert) have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability (rives2021biological).
 For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes.
 (3) Inefficiency for Long Sequences.
-Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length alquraishi2019proteinnet .
+Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length (alquraishi2019proteinnet).
 This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
+Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM), separately proposed.
 Figure 2: Illustration of Life-Code Framework, which contains three pipelines.
@@
 (b) Pre-training of the Tokenizer and the hybrid Encoder to model the contextual information and the central dogma rules.
-(c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning (SFT).
+(c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning.
 This paper introduces Life-Code, as shown in Figure 2, a unifying framework that addresses these limitations through multi-omics unification, bio-interpretable designs, and long-sequence efficiency.
@@
 Linear or sparse approximations (e.g., Mamba) reduce complexity but often sacrifice representational fidelity for complex biological contexts.
-Hence, the recent efficient hybrid strategy yang2025deltanet ; yang2024gateddelta ; iclr2024MogaNet ; icml2024chela ; ijcai2024longvq should be considered.
+Hence, the recent efficient hybrid strategy (yang2025deltanet; yang2024gateddelta; iclr2024MogaNet; icml2024chela; ijcai2024longvq) should be considered.
 We now formally describe how we unify DNA, RNA, and protein sources into nucleotides.
@@
 Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample.
-As for the CDS and its corresponding amino acids, we apply the packing strategy Warner2024ModernBERT to sample with several CDSs.
+As for the CDS and its corresponding amino acids, we apply the packing strategy (Warner2024ModernBERT) to sample with several CDSs.
 Data sampling.
-Following GenBank benson2012genbank curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones.
+Following GenBank (benson2012genbank) curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones.
 For CDS–Amino Acid pairs, we extract annotated CDS records and their corresponding protein sequences, then pack multiple short segments into a single training sample with special separators.
@@
 By balancing truncation/padding for DNA sequences and packing for CDS segments, we ensure that each batch contains a diverse mixture of long-range genomic context and coding information, thereby maximizing the model’s exposure to various biological signals.
-Figure 4: Life-Code Tokenizer.
-To model the interactions among DNA, RNA, and Amino Acids, our tokenizer takes nucleotide acids (4 words) as the inputs, then takes codons as the latent vocabulary (64 words), and translates them to Amino Acids (20 words) as the outputs, which could be pre-trained by (a) nucleotide masked modeling and (b) CDS to Amino Acid translation.
 2.3 Life-Code Tokenizer Training
@@
 Consider a raw DNA segment of length $3L$.
-We first apply a 1-d convolution (kernel size $=3$) together with an “unfold” operation (stride $=3$) to transform each consecutive triplet $(x_{3j+1},x_{3j+2},x_{3j+3})$ into a codon-like representation:
-$\mathbf{e}_{j}\;=\;\mathrm{Conv1d}(\,x_{3j+1:3j+3}\,).$
+We first apply a 1-d convolution (kernel size =3=3) together with an “unfold” operation (stride =3=3) to transform each consecutive triplet $(x_{3j+1},x_{3j+2},x_{3j+3})$ into a codon-like representation:
+$\mathbf{e}_{j}\;=\;\mathrm{Conv1d}(\,\mathbf{x}_{3j+1:3j+3}\,).$
 This step effectively condenses triplets—codons—into a more compact embedding sequence $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{L}\}\in\mathbb{R}^{L\times 3d}$.
@@
 Formally, if $\widehat{\mathbf{x}}_{3j+1:3j+3}$ is the predicted codon for position $j$, the loss is:
-$\mathcal{L}_{\mathrm{MLM}}\;=\;\sum_{j=1}^{L}-\log p\bigl{(}\mathbf{x}_{3j+1:3j+3}\,\big{|}\,\widetilde{\mathbf{x}}\bigr{)},$ (2)
+[…]
 where $\widetilde{\mathbf{x}}$ is the masked sequence.
@@
 By training on paired CDS–amino-acid data, the tokenizer acquires biologically grounded representations, effectively bridging codon-level features to protein-level semantics.
+Figure 4: Life-Code Tokenizer.
+To model the interactions among DNA, RNA, and Amino Acids, our tokenizer takes nucleotide acids (4 words) as the inputs, then takes codons as the latent vocabulary (64 words), and translates them to Amino Acids (20 words) as the outputs, which could be pre-trained by (a) nucleotide masked modeling and (b) CDS to Amino Acid translation.
 2.4 Bi-Directional Hybrid Model for Long Sequence
@@
 Bi-directional input.
-To use the double-helix property as Caduceus (icml2024caduceus, ), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
+To use the double-helix property as Caduceus (icml2024caduceus), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
 $(\mathbf{E}^{(+)},\ \mathbf{E}^{(-)})=\mathrm{Split}(\mathbf{E}),$
@@
 Finally, we concatenate or fuse the representations:
-[…]
+$\mathbf{H}=[\mathbf{H}^{(+)},\ \mathbf{H}^{(-)}],$
 where $\mathbf{H}\in\mathbb{R}^{D}$ encodes both the forward and reverse complement features.
+Figure 5: Empirical Analysis of Life-Code Tokenizer.
+Left: Codon Usage Bias in Life-Code Tokenizer across four representative species—E.
+coli, S.
+cerevisiae (yeast), D.
+melanogaster (fruit fly), and H.
+sapiens (human)—illustrating variations in codon frequency (%) for amino acids.
+The pastel strips highlight codons belonging to the same amino acid group.
+Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons form a distinct region.
 Hybrid Token Mixer.
-Inspired by recent advances in EVO-style architectures nguyen2024evo and Gated Delta Networks (GDN) yang2024gateddelta , we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA).
+Inspired by recent advances in EVO-style architectures (nguyen2024evo) and Gated Delta Networks (GDN) (yang2024gateddelta), we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA).
 Concretely, for every twelve encoder layers, eleven employ the GDN update rule—offering linear or near-linear complexity via kernel-based reordering—while one layer applies MHSA.
@@
 For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues).
-Inspired by the DINO framework (oquab2024dinov2, ), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (lin2022ESM2, ).
+Inspired by the DINO framework (oquab2024dinov2), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (lin2022ESM2).
 Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 4D\rightarrow D^{\prime}$.
@@
 $\mathcal{L}_{\mathrm{KD}}\;=\;\sum_{j=1}^{L}-P_{\theta^{\prime}}\bigl(\mathbf{h}_{j}^{\mathrm{PLM}}\bigr)^{\top}\log P_{\theta}\bigl(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}})\bigr).$ (4)
-We combine these three losses using a log-sum multi-task strategy (lin2023dualmtl, ) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
-$\log\bigl{(}\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr{)},$ (5)
+We combine these three losses using a log-sum multi-task strategy (lin2023dualmtl) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
+$\hskip-10.00002pt\log\bigl(\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr)\;+\;\log\bigl(\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr)\;+\;\log\bigl(\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr),$ (5)
 where $\epsilon$ is a small constant.
@@
 Notably, we also adopt this strategy when training the tokenizer.
+Figure 6: Training Efficiency on a Single A100 GPU.
+We measure throughput (K t/s) at four sequence lengths–batch size settings.
+DNABERT-2 slows significantly at higher lengths and reaches an out-of-memory (OOM) at $16\text{K}\times 2$.
 Figure 7: Illustration of Life-Code Encoder pre-training pipeline.
-We model the long sequence by the hybrid encoder as a mixture of RNN blocks (DeltaNet) and self-attention blocks.
-With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma (with span masks), and (c) the knowledge distillation (KD) with pre-trained protein LM.
+We model the long sequence by the hybrid encoder as a mixture of DeltaNet and Attention blocks.
+With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma, and (c) the knowledge distillation (KD) with pre-trained protein LM.
 Note that red lines indicates the learning objectives.
-Figure 6: Empirical Analysis of Life-Code Tokenizer.
-Left: Codon Usage Bias in Life-Code Tokenizer across four representative species—E.
-coli, S.
-cerevisiae (yeast), D.
-melanogaster (fruit fly), and H.
-sapiens (human)—illustrating variations in codon frequency (%) for amino acids.
-The pastel strips highlight codons belonging to the same amino acid group.
-Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons (black crosses) form a distinct region.
-These views jointly demonstrate that Life-Code encodes both species-dependent codon usage preferences and biologically meaningful relationships.
 2.5 Empirical Consequences
@@
 This structural arrangement confirms that the model internalizes codon–amino acid mappings, reflecting both frequency patterns and higher-level attributes like polarity or aromaticity.
-Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals—ranging from species-specific usage preferences.
-Figure 7: Training Efficiency on a Single A100 GPU.
-We measure throughput (K t/s) at four sequence lengths–batch size settings.
-DNABERT-2 slows significantly at higher lengths and reaches an out-of-memory (OOM) at $16\text{K}\times 2$.
+Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals.
 Training efficiency
@@
 Notably, Life-Code achieves the best scalability, retaining above 40K t/s even at $8\text{K}\times 4$.
-These results confirm that Life-Code’s efficient architecture can accommodate extensive input lengths with minimal degradation in training speed, making it well-suited for large-scale genomic modeling.
+These results confirm that Life-Code’s efficient architecture can accommodate extensive input lengths with minimal degradation in training speed.
 3 Experiments
@@
 Similarly, we also design the DNA De-Tokenizer and Amino Acid Translator for tokenizer pre-training.
-As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (iclr2019AdamW, ) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2.
-As shown in Table A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases, as described in Figure A1, Appendix A, and  A.2.
+As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (iclr2019AdamW) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2.
+As shown in Table A1 and  A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases.
 Life-Code Encoder.
-Similar to the BERT-Large architecture (devlin2019bert, ), the Life-Code encoder has 24 layers in total with an embedding dim of 1024/1280 and 340/500M parameters.
-We enhance the model with three-fold designs: (i) Mixture of DeltaNet (yang2025deltanet, ) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block.
-(ii) The LLaMA-like macro design (Touvron2023LLaMA, ) with RMSNorm (Zhang2019RMSNorm, ), Layer Scale (iccv2021CaiT, ), Rotary Position Embedding (RoPE) (Su2021RoFormer, ), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences.
-(iii) During pre-training, we apply the packing strategy (Warner2024ModernBERT, ) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences.
+Similar to the BERT-Large architecture (devlin2019bert), the Life-Code encoder has 24 layers in total with an embedding dim of 1024/1280 and 340/500M parameters.
+We enhance the model with three-fold designs: (i) Mixture of DeltaNet (yang2025deltanet) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block.
+(ii) The LLaMA-like macro design (Touvron2023LLaMA) with RMSNorm (Zhang2019RMSNorm), Layer Scale (iccv2021CaiT), Rotary Position Embedding (RoPE) (Su2021RoFormer), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences.
+(iii) During pre-training, we apply the packing strategy (Warner2024ModernBERT) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences.
 With three pre-training tasks in Eq.
@@
 We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation.
-As for KD from a pre-trained Protein LM, we adopted ESM2-650M (lin2022ESM2, ) and a protein decoder with an output dimension of 1280.
+As for KD from a pre-trained Protein LM, we adopted ESM2-650M (lin2022ESM2) and a protein decoder with an output dimension of 1280.
 During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations.
@@
 View Appendix A.2 for more details.
-Table 1: Comparison on Genomic Benchmarks.
-Top-1 accuracy (%) averaged over several similar tasks is reported for popular DNA foundation models with SFT evaluation.
-The best and the second best results are marked as the bold and underlined types.
 Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
@@
 Regulatory Elements (3 tasks) 88.89 83.42 86.33 87.89 89.05 85.97 87.62 90.57 87.30 91.30
-Average 87.07 85.02 87.30 87.94 89.26 85.89 87.69 89.12 86.95 90.79
-Table 2: Comparison on GUE Benchmark.
-MCC (%) averaged across several sub-tasks is reported with SFT evaluation for seven primary tasks (24 sub-tasks in total).
+Average (8 tasks) 87.07 85.02 87.30 87.94 89.26 85.89 87.69 89.12 86.95 90.79
+Table 1: Comparison on Genomic Benchmarks.
+Top-1 accuracy (%) averaged over several similar tasks is reported for popular DNA foundation models with SFT evaluation.
 The best and the second best results are marked as the bold and underlined types.
@@
 Epigenetic Marks Prediction (10) 58.94 49.08 58.06 55.98 58.39 57.95 67.29 68.91 67.85
-Human TF Detection (3) -- 64.17 63.34 70.11 -- 70.56 -- -- 71.34
-Mouse TF Detection (3) -- 56.43 67.02 67.99 -- 69.80 -- -- 71.76
-Core Promoter Detection (3) -- 71.81 71.63 70.53 -- 73.37 -- -- 74.53
-Promoter Detection (3) -- 81.69 88.15 84.21 -- 86.58 -- -- 89.28
-Splice Site Reconstructed (1) -- 84.07 89.35 84.99 -- 89.53 -- -- 90.24
-Virus Covid Classification (1) -- 55.50 73.04 71.02 -- 74.32 -- -- 74.88
+Human TF Detection (3) - 64.17 63.34 70.11 - 70.56 - - 71.34
+Mouse TF Detection (3) - 56.43 67.02 67.99 - 69.80 - - 71.76
+Core Promoter Detection (3) - 71.81 71.63 70.53 - 73.37 - - 74.53
+Promoter Detection (3) - 81.69 88.15 84.21 - 86.58 - - 89.28
+Splice Site Reconstructed (1) - 84.07 89.35 84.99 - 89.53 - - 90.24
+Virus Covid Classification (1) - 55.50 73.04 71.02 - 74.32 - - 74.88
 Average (24 tasks) 58.94 60.53 67.23 66.43 58.39 68.51 67.29 68.91 73.51
+Table 2: Comparison on GUE Benchmark.
+Matthews Correlation Coefficient (MCC) (%) averaged across several sub-tasks is reported with SFT evaluation.
+The best and the second best results are marked as the bold and underlined types.
 3.2 Comparison on Genomic Benchmarks
-As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna, ; iclr2024dnabert2, ), following benchmark evaluations for fair comparison.
-We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert, ), DNABERT2 (iclr2024dnabert2, ), GENA-LM (fishman2023GenaLM, ), Nucleotide Transformer (NT) (NM2023NucleotideTrans, ), with different tokenizers like VQDNA (icml2024vqdna, ) and MxDNA (nips2024MXDNA, ); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna, ) and Caduceus (icml2024caduceus, ); (c) CNN model ConvNova (iclr2025convnova, ).
+As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna; iclr2024dnabert2), following benchmark evaluations for fair comparison.
+We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert), DNABERT2 (iclr2024dnabert2), GENA-LM (fishman2023GenaLM), Nucleotide Transformer (NT) (NM2023NucleotideTrans), with different tokenizers like VQDNA (icml2024vqdna) and MxDNA (nips2024MXDNA); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna) and Caduceus (icml2024caduceus); (c) CNN model ConvNova (iclr2025convnova).
 Genomic Benchmarks.
-We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark, ), which we adopted results from GenBench implementations (liu2024genbench, ), and all compared DNA models are fully SFT and reported the top-1 accuracy.
-Table 1 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages.
-GUE Benchmarks.
-We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2, ), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification.
-All results are adopted from DNABERT2 or each paper with their official SFT setups, and the Matthews Correlation Coefficient (MCC) is reported as the metric.
-3.3 Comparison on RNA Splicing Site Prediction
-Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks.
-We consider two typical datasets with the SFT evaluation protocol.
-As for the SpliceAI dataset (JAGANATHAN2019SpliceAI, ), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma.
-As for Spliceator dataset (BMC2021spliceator, ), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP, ), RINALMo (Penic2024RiNALMo, )) and RNA language model, RNA-FM (shen2024rnafm, ), with RNA data of multi-species.
-As shown in Table 4, our Life-Code achieves the best F1 scores with the specified baselines.
-Table 3: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets.
+We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark), which we adopted results from GenBench implementations (liu2024genbench), and all compared DNA models are fully SFT and reported the top-1 accuracy.
+Table 1 demonstrates that Life-Code surpasses previous models since its hybrid designs combine the advantages.
 Method GenSLM NT-2500M EVO Life-Code
@@
 Human 6.9 4.7 11.1 23.8
-Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
+Table 3: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets.
+GUE Benchmarks.
+We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification.
+All results are adopted from DNABERT2 or each paper with their official SFT setups.
+3.3 Comparison on RNA Splicing Site Prediction
+Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks.
+We consider two typical datasets with the SFT evaluation protocol.
+As for the SpliceAI dataset (JAGANATHAN2019SpliceAI), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma.
+As for Spliceator dataset (BMC2021spliceator), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP), RINALMo (Penic2024RiNALMo)) and RNA language model, RNA-FM (shen2024rnafm), with RNA data of multi-species.
+As shown in Table 4, our Life-Code achieves the best F1 scores.
 Dataset SpliceAI DNABERT2 NT Caduceus Life-Code
@@
 Spliceator 91.0 91.9 92.1 94.8 95.0
+Table 4: Splicing Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
 3.4 Comparison on Protein Fitness Prediction
-Following EVO (nguyen2024evo, ), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction.
-Deep Mutational Scanning (DMS) datasets (icml2022Tranception, ; nguyen2024evo, ) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy.
+Following EVO (nguyen2024evo), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction.
+Deep Mutational Scanning (DMS) datasets (icml2022Tranception; nguyen2024evo) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy.
 For both bacterial and human proteins, we compute Spearman’s rank correlation (SRCC) between model scores and experimental measurements (Table 3).
-The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms, ) and the 7-B-parameter EVO model—despite using fewer parameters.
-We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without any task-specific fine-tuning.
-Table 5: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation.
+The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms) and the 7-B-parameter EVO model—despite using fewer parameters.
+We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without fine-tuning.
+3.5 Comparison on Multi-omics Tasks
+We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo) and LucaOne (he2024lucaone)) with multi-omics tasks.
+Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between non-coding RNAs (ncRNAs) and proteins.
+As shown in Table 5, Life-Code can accept cDNA and Amino Acid sequences in a unified input format and achieve a competitive result, comparable to the supervised LucaOne, which requires separate modules for each input.
+Then, we evaluate the modeling of transcription and translation processes using binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy compared to existing models.
+We can see that the modalities enhance each other.
 Method ncRNA-Protein Central Dogma
 DNABERT2+ESM2 93.32 73.31
-EVO -- 75.45
+EVO - 75.45
 LucaOne 93.80 84.84
 Life-Code 94.12 86.51
-Table 6: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory on the single A100-80G.
+Table 5: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation.
 Method # Params 1k 20k 32k 100k 250k
@@
 Life-Code 500M 62.8 88.1 99.3 99.4 99.5
-Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss; 3:1 exceeds memory limits.
-3.5 Comparison on Multi-omics Tasks
-We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo, ) and LucaOne (he2024lucaone, )) with multi-omics tasks.
-Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins.
-As shown in Table 5, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve a competitive result as the supervised LucaOne, which should process two inputs with separate modules.
-Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy over existing models.
-3.6 Empirical Analysis
+Table 6: Comparison on Species classification. Top-1 accuracy (%) for 5-way classification with scaling up the sequence length, where OOM denotes out-of-memory.
+Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss.
+3.6 Empirical Analysis of Hybrid Design
 Figure 8 shows that a pure RNN stack (12:0) is fastest ( $\sim$45 K token/s) but sacrifices 2–3 pp of accuracy.
@@
 A 5:1 mix boosts accuracy further with a modest extra cost, whereas 3:1 exhausts memory.
-Thus, adding a small fraction of MHSA markedly improves downstream accuracy without a large computational penalty.
-Table 7: Ablation Study of Life-Code tokenizer, encoder, and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
+Thus, adding a small fraction of MHSA markedly improves downstream efficiency.
 Tokenizer Encoder PT Tasks DNA Protein CD
-Linear (1-mer) MLM 58.3 -- 28.7
+Linear (1-mer) MLM 58.3 - 28.7
 GDN (codon) MLM+Trans. 84.5 6.4 63.2
@@
 GDN (codon) Hybrid MLM+Trans.+KD 97.0 34.1 85.6
+Table 7: Ablation Study of Life-Code and pre-training tasks with DNA, protein, and central dogma (CD) tasks. Note that the blue background denotes the selected setups.
 3.7 Ablation Study
@@
 DNA Foundation Models.
-Recent DNA foundation models have evolved swiftly from DNABERT ji2021dnabert , which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 zhou2023dnabert2 and the large-scale Nucleotide Transformer NM2023NucleotideTrans .
-To tame quadratic attention on longer sequences, HyenaDNA nguyen2024hyenadna replaces full self-attention with efficient hierarchical kernels.
-Other variants embed biological priors—Caduceus schiff2024caduceus for long-range structure and MxDNA nips2024MXDNA for adaptive tokenization—illustrating a broad shift toward specialized, scalable architectures for genomics.
+Recent DNA foundation models have evolved swiftly from DNABERT (ji2021dnabert), which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 (zhou2023dnabert2) and the large-scale Nucleotide Transformer (NM2023NucleotideTrans).
+To tame quadratic attention on longer sequences, HyenaDNA (nguyen2024hyenadna) replaces full self-attention with efficient hierarchical kernels.
+Other variants embed biological priors—Caduceus (schiff2024caduceus) for long-range structure and MxDNA (nips2024MXDNA) for adaptive tokenizer.
 Protein and RNA Advances.
-AlphaFold jumper2021alphafold ; jumper2021alphafold2 ignited the protein-LM wave, inspiring models like ESM-2 lin2022ESM2 .
-RNA efforts followed, e.g., RNA-FM shen2024rnafm for 3D structure prediction.
-While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals prakash2024bridging ; bioinfo2024genomic ; BioInfo2023NTA ; nmil2024CaLM , delivering notable downstream gains Ren2024COMET and pointing toward unified multi-omics models.
+AlphaFold (jumper2021alphafold; jumper2021alphafold2) ignited the protein-LM wave, inspiring models like ESM-2 (lin2022ESM2).
+RNA efforts followed, e.g., RNA-FM (shen2024rnafm) for 3D structure prediction.
+While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals (prakash2024bridging; bioinfo2024genomic; BioInfo2023NTA; nmil2024CaLM), delivering notable downstream gains (Ren2024COMET) and toward unified models.
 Multi-Omics Modeling.
 Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks.
-CD-GPT zhu2024CDGPT explicitly connects DNA, RNA, and proteins through the central dogma, while BSM xiang2024BSM highlights the potential for small but effective models covering genes and proteins simultaneously.
-EVO nguyen2024evo similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach.
-Closely related is LucaOne he2024lucaone , which advocates a unified nucleic acid and protein language for biological representation.
+CD-GPT (zhu2024CDGPT) explicitly connects DNA, RNA, and proteins through the central dogma, while BSM (xiang2024BSM) highlights the potential for small but effective models covering genes and proteins simultaneously.
+EVO (nguyen2024evo) similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach.
+Closely related is LucaOne (he2024lucaone), which advocates a unified nucleic acid and protein language for biological representation.
 5 Conclusion and Limitations
@@
 Despite these challenges, we believe Life-Code offers a promising step toward comprehensive multi-omics data.
-Appendix
+Appendix of Life-Code
   • In Appendix A, we provide implementation details of pre-training datasets, network architectures, and training schemes of pre-training and fine-tuning stages with hyper-parameter settings.
@@
 Figure A1: Data Collection Pipeline.
-We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA pairing datasets.
-Table A1: Configuration of pre-training datasets.
-As for the DNA dataset, we collect DNA/RNA sequences from the NCBI RefSeq database.
-As for the DNA-AA pairing dataset, we collect the cDNA (coding sequences) and its corresponding Amino Acids (or using translation) from the GenBank database and also collect Amino Acids and their corresponding cDNA sequences using reverse translation from the UniRef50 database.
+We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build our DNA-AA pairing datasets.
 Dateset Data Type Seq Count Data Source
@@
 RNA 6,463,852 NCBI RefSeq
-cDNA-AA pair 16,786,593 GenBank
+cDNA-AA pair 49,453,295 GenBank
 DNA-AA AA-cDNA pair 17,245,138 UniRef50
-mRNA-AA pair 3.908,074 GenBank
+mRNA-AA pair 25,468,310 GenBank
+Table A1: Configuration of pre-training datasets.
+As for the DNA dataset, we collect DNA/RNA sequences from the NCBI RefSeq database.
+As for the DNA-AA pairing dataset, we collect the cDNA (coding sequences) and their corresponding Amino Acids (or using translation) from the GenBank database, and also collect Amino Acids and their corresponding cDNA sequences using reverse translation from the UniRef50 database.
 Appendix A Implementation Details
 A.1 Pre-training Dataset
-We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1.
-As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer [8] and DNABERT2 [64].
-As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma.
-We also collect some Amino Acids in the UniRef50 database following LucaOne [13] and obtain their corresponding cDNA by reverse translation with online tools.
-We provide detailed information for used datasets in Table A1.
+We collect two datasets for pre-training Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, as shown in Figure A1.
+As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi˙species˙genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT˙2 provided by Nucleotide Transformer (NM2023NucleotideTrans) and DNABERT2 (iclr2024dnabert2).
+It includes around 135 species in 7 categories with 35.4B nucleotide bases in total.
+As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and their corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma.
+We also collect some Amino Acids in the UniRef50 database following LucaOne (he2024lucaone) and obtain their corresponding cDNA by reverse translation with online tools to build cDNA-AA pairs.
+It contains 92M sequences in total.
+We provide detailed information for the used datasets in Table A1.
 A.2 Life-Code Tokenizer
@@
 Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”).
-It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS).
-The pre-trained protein language model employs the amino acid vocabulary of ESM-2 [26].
-Table A2: Configuration of the network designs and pre-training settings for Life-Code models.
-GDN blocks denote the Gated DeltaNet block with linear complexity, while Attention blocks denote the self-attention block with FLASH-Attention implementation.
+It can only be applied when the length of an input sequence is a multiple of 3, i.e., the cDNA of amino acids or matured mRNA (CDS).
+The pre-trained protein language model employs the amino acid vocabulary of ESM-2 (lin2022ESM2).
 Configuration Tokenizer Encoder
@@
 Attention blocks 0 2
-Attention heads 0 24
-Parameters 8M 350/500M
+Attention heads 0 16
+Parameters 8M 305/500M
 Optimizer AdamW
@@
 Gradient clipping 1.0
+Table A2: Configuration of the network designs and pre-training settings for Life-Code models.
+GDN blocks denote the Gated DeltaNet (yang2024gateddelta) block with linear complexity, while Attention blocks denote the self-attention block with FLASH-Attention implementation (Dao2022FlashAttention).
 Tokenizer Network.
@@
 Encoder Architecture.
-As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet [60]) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 [32], which could utilize the complementary properties of GDN and MHSA while maintaining efficiency.
-Notice that the DeltaNet variants [59] introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block.
-(2) The model macro design employs pre-norm [55] with RMSNorm [61], Layer Scale [52], Rotary Position Embedding (RoPE) [51], SwiGLU [53], and FlashAttention implementations to facilitate training large-scale models stably with long sequences.
-(3) During pre-training, we apply the packing strategy [57] to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
+As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet (yang2025deltanet)) and multi-head self-attention (MHSA) blocks with the head dim of 64 as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 (MiniMax2025MiniMax01), which could utilize the complementary properties of GDN and MHSA while maintaining efficiency.
+Notice that the DeltaNet variants (yang2024gateddelta) introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block.
+(2) The model macro design employs pre-norm (acl2019PreNorm) with RMSNorm (Zhang2019RMSNorm), Layer Scale (iccv2021CaiT), Rotary Position Embedding (RoPE) (Su2021RoFormer), SwiGLU (Touvron2023LLaMA), and FlashAttention implementations to facilitate training large-scale models stably with long sequences.
+(3) During pre-training, we apply the packing strategy (Warner2024ModernBERT) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
 Pre-training Settings.
@@
 We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation.
-As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D[26] and a protein decoder with the output dimension of 1280.
+As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D(lin2022ESM2) and a protein decoder with the output dimension of 1280.
 During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations.
@@
 In most cases, we apply Supervised Fine-tuning (SFT) to transfer pre-trained models to downstream tasks.
-Following [36, 64], adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy [15] is employed to parameter-efficiently fine-tuning the models by AdamW optimizer with a batch size of 32.
-For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate $1\times 10^{-5}$, $5\times 10^{-5}$, $1\times 10^{-4}$}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench [28].
+Following (nips2024hyenadna; iclr2024dnabert2), adding the decoder head (e.g., an MLP head) to a specific downstream task, the linear attention (RNN) or self-attention blocks in the pre-trained encoder models are frozen, while Low-Rank Adaptation (LoRA) strategy (Hu2021LoRA) is employed to parameter-efficiently fine-tuning the models by the AdamW optimizer with a batch size of 32.
+For each task, if the benchmark and models have provided hyper-parameters, we follow the official settings, or we choose the best combinations of the basic learning rate $1\times 10^{-5}$, $5\times 10^{-5}$, $1\times 10^{-4}$}, the weight decay {0, 0.01}, the LoRA rank {4, 8, 16, 24, 48}, the LoRA alpha {8, 16, 24, 48, 96}, and the total fine-tuning epoch {5, 10} on the validation set following the GUE benchmark and GenBench (liu2024genbench).
 Note that the maximum input length will be determined for different tasks since the sequence lengths of downstream tasks vary widely.
 We report the averaged results over three runs with the optimal settings.
-Table A3: Full Results on Genomic Benchmarks.
-Top-1 accuracy (%) averaged across three trials is reported for the latest DNA foundation models, where the best and the second best results are marked as the bold and underlined types.
 Method HyenaDNA DNABERT DNABERT2 GENA-LM NT-500M Caduceus-16 VQDNA MxDNA ConvNova Life-Code
@@
 Human NonTATA Promoters 94.45 87.13 95.24 96.60 92.95 88.85 95.37 96.56 95.30 96.65
-Table A4: Full Results on GUE benchmark.
-MCC (%) is reported for Epigenetic Marks Prediction, Human Transcription Factor (TF) Prediction, Mouse Transcription Factor Prediction, Core Promoter Detection, Promoter Detection, Splice Site Reconstructed, and Covid Variants Classification (Virus Covid).
-The best and the second best results are marked as the bold and underlined types.
+Table A3: Full Results on Genomic Benchmarks.
+Top-1 accuracy (%) averaged across three trials is reported for the latest DNA foundation models, where the best and the second best results are marked as the bold and underlined types.
 Method HyenaDNA DNABERT NT-2500M-multi DNABERT2 Caduceus-PS VQDNA MxDNA ConvNova Life-Code
 # Params (M) 6.6 86 2537 117 1.9 93 100 1.7 350
-Human TF-0 -- 66.84 66.64 71.99 -- 72.48 -- -- 71.58
-Human TF-1 -- 70.14 70.28 76.06 -- 76.43 -- -- 75.92
-Human TF-2 -- 61.03 58.72 66.52 -- 66.85 -- -- 70.63
-Human TF-3 -- 51.89 51.65 58.54 -- 58.92 -- -- 58.74
-Human TF-4 -- 70.97 69.43 77.43 -- 78.10 -- -- 79.45
-Mouse TF-0 -- 44.42 63.31 56.76 -- 58.34 -- -- 64.10
-Mouse TF-1 -- 78.94 83.76 84.77 -- 85.81 -- -- 86.51
-Mouse TF-2 -- 71.44 71.52 79.32 -- 80.39 -- -- 80.49
-Mouse TF-3 -- 44.89 69.44 66.47 -- 69.72 -- -- 71.25
-Mouse TF-4 -- 42.48 47.07 52.66 -- 54.73 -- -- 55.46
-Core Promoter (all) -- 68.90 70.33 69.37 -- 71.02 -- -- 70.69
-Core Promoter (no TATA) -- 70.47 71.58 68.04 -- 70.58 -- -- 71.05
-Core Promoter (TATA) -- 76.06 72.97 74.17 -- 78.50 -- -- 78.78
-Promoter (all) -- 90.48 91.01 86.77 -- 90.75 -- -- 91.33
-Promoter (no TATA) -- 93.05 94.00 94.27 -- 94.48 -- -- 95.03
-Promoter (TATA) -- 61.56 79.43 71.59 -- 74.52 -- -- 78.97
-Splice Reconstructed -- 84.07 89.35 84.99 -- 89.53 -- -- 89.76
+Human TF-0 - 66.84 66.64 71.99 - 72.48 - - 71.58
+Human TF-1 - 70.14 70.28 76.06 - 76.43 - - 75.92
+Human TF-2 - 61.03 58.72 66.52 - 66.85 - - 70.63
+Human TF-3 - 51.89 51.65 58.54 - 58.92 - - 58.74
+Human TF-4 - 70.97 69.43 77.43 - 78.10 - - 79.45
+Mouse TF-0 - 44.42 63.31 56.76 - 58.34 - - 64.10
+Mouse TF-1 - 78.94 83.76 84.77 - 85.81 - - 86.51
+Mouse TF-2 - 71.44 71.52 79.32 - 80.39 - - 80.49
+Mouse TF-3 - 44.89 69.44 66.47 - 69.72 - - 71.25
+Mouse TF-4 - 42.48 47.07 52.66 - 54.73 - - 55.46
+Core Promoter (all) - 68.90 70.33 69.37 - 71.02 - - 70.69
+Core Promoter (no TATA) - 70.47 71.58 68.04 - 70.58 - - 71.05
+Core Promoter (TATA) - 76.06 72.97 74.17 - 78.50 - - 78.78
+Promoter (all) - 90.48 91.01 86.77 - 90.75 - - 91.33
+Promoter (no TATA) - 93.05 94.00 94.27 - 94.48 - - 95.03
+Promoter (TATA) - 61.56 79.43 71.59 - 74.52 - - 78.97
+Splice Reconstructed - 84.07 89.35 84.99 - 89.53 - - 89.76
 H3 78.14 73.10 78.77 78.27 77.90 79.21 82.14 81.50 81.28
@@
 H4ac 54.15 37.24 49.13 50.43 52.50 50.69 65.14 66.10 61.37
-Virus Covid Classification -- 55.50 73.04 71.02 -- 74.32 -- -- 73.82
+Virus Covid Classification - 55.50 73.04 71.02 - 74.32 - - 73.82
+Table A4: Full Results on GUE benchmark.
+MCC (%) is reported for Epigenetic Marks Prediction, Human Transcription Factor (TF) Prediction, Mouse Transcription Factor Prediction, Core Promoter Detection, Promoter Detection, Splice Site Reconstructed, and Covid Variants Classification (Virus Covid).
+The best and the second best results are marked as the bold and underlined types.
 Appendix B Downstream Task Settings and Extensive Comparison Results
 B.1 DNA Tasks with Genomics Benchmark
-As proposed by [12], three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark.
+As proposed by (BMC2023genomicbenchmark), three groups of basic genomic tasks are collected as binary classification with top-1 accuracy in the Genomics Benchmark.
 As for the enhancer prediction, three datasets are provided for identifying enhancer regions in the mouse or human genome.
@@
 As for the regulatory elements classification, three datasets are used for classifying sequences as regulatory regions based on Ensembl annotations, identifying open chromatin regions, or identifying non-TATA promoter regions in the human genome.
-We utilize the fully reproduced results of various DNA models in GenBench [28].
+We utilize the fully reproduced results of various DNA models in GenBench (liu2024genbench).
 B.2 DNA Tasks with GUE Benchmark
-As proposed by DNABERT2 [64], the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric.
+As proposed by DNABERT2 (iclr2024dnabert2), the GUE benchmark contains 24 datasets of 7 practical biological genome analysis tasks for 4 different species using Matthews Correlation Coefficient (MCC) as the evaluation metric.
 To comprehensively evaluate the genome foundation models in modeling variable-length sequences, tasks with input lengths ranging from 70 to 1000 are selected.
@@
 Accurate detection aids in understanding gene regulation and disease mechanisms.
-The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) [10].
+The dataset includes TATA and non-TATA promoters, with sequences -249 to +50 bp around the Transition Start Site (TSS) from Eukaryotic Promoter Database (EPDnew) (dreos2013epdpromoter).
 Meanwhile, we construct the non-promoter class with equal-sized randomly selected sequences outside of promoter regions but with TATA motif (TATA non-promoters) or randomly substituted sequences (non-TATA, non-promoters).
@@
 This task predicts human transcription factor binding sites (TFBS), crucial for gene expression regulation.
-Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) [6] are collected via the UCSC genome browser.
+Data from 690 ENCODE ChIP-seq experiments (161 TF binding profiles in 91 cell lines) (mouse2012encyclopedia) are collected via the UCSC genome browser.
 TFBS sequences are 101-bp regions around peaks, while non-TFBS sequences match in length and GC content.
@@
 This prediction is crucial to understanding protein diversity and the implications of aberrant splicing in genetic disorders.
-The dataset [46] consists of 400-bp-long sequences extracted from Ensembl GRCh38 human reference genome.
-As suggested by [17], existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality.
+The dataset (BMC2021spliceator) consists of 400-bp-long sequences extracted from the Ensembl GRCh38 human reference genome.
+As suggested by (BioInfo2021dnabert), existing models can achieve almost perfect performance on the original dataset, containing 10,000 splice donors, acceptors, and non-splice site sequences, which is overly optimistic about detecting non-canonical sites in reality.
 As such, we reconstruct the dataset by iteratively adding adversarial examples (unseen false positive predictions in the hold-out set) in order to make this task more challenging.
 Transcription Factor Binding Site Prediction (Mouse).
-This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) [7] from the UCSC genome browser.
+This task predicts mouse transcription factor binding sites using mouse ENCODE ChIP-seq data (n=78) (mouse2012ChIPseq) from the UCSC genome browser.
 Negative examples are created by di-nucleotide shuffling.
@@
 Precise prediction of these marks aids in elucidating the role of epigenetics in yeast.
-We download the 10 datasets from http://www.jaist.ac.jp/~tran/nucleosome/members.htm and randomly split each dataset into training, validation, and test sets.
+We download the 10 datasets from http://www.jaist.ac.jp/˜tran/nucleosome/members.htm and randomly split each dataset into training, validation, and test sets.
 Covid Variant Prediction (Virus).
 This task aims to predict the variant type of the SARS_CoV_2 virus based on 1000-length genome sequences.
-We download the genomes from the EpiCoV database [21] of the Global Initiative on Sharing Avian Influenza Data (GISAID).
+We download the genomes from the EpiCoV database (khare202SARS_CoV_2) of the Global Initiative on Sharing Avian Influenza Data (GISAID).
 We consider 9 types of SARS_CoV_2 variants, including Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda and Zeta.
-Table A5: Full Results of mRNA Splicing Site Prediction.
-With two splicing datasets proposed by SpliceAI and Spliceator, the top-1 AUC score or F1 score is reported with DNA models or RNA models, respectively.
-The best and the second best results are marked as the bold and underlined types.
 Dataset SpliceAI DNABERT2 NT GENA-LM Caduceus Life-Code
@@
 Fly 91.0 84.2 91.4 91.9 95.8 94.7
+Table A5: Full Results of mRNA Splicing Site Prediction.
+With two splicing datasets proposed by SpliceAI and Spliceator, the top-1 AUC score or F1 score is reported with DNA models or RNA models, respectively.
+The best and the second best results are marked as the bold and underlined types.
 B.3 mRNA Splicing Tasks
-Following [64, 49], we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression.
+Following (iclr2024dnabert2; shen2024rnafm), we evaluate pre-mRNA Splicing Site Prediction as the RNA task, which is a crucial process in eukaryotic gene expression.
 During splicing, introns are removed from precursor messenger RNAs (pre-mRNAs), and exons are joined together to form mature mRNAs.
@@
 Identifying splice sites—the donor sites at the 5’ end of introns and the acceptor sites at the 3’ end—is vital for accurately predicting gene structure and location.
-Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset [56] used in the GUE benchmark.
+Concretely, we regard this task as binary classification of RNA splicing site prediction specifically for acceptor sites and consider two splicing datasets in addition to the Splicefinder dataset (wang2019splicefinder) used in the GUE benchmark.
 Spliceator dataset.
-This dataset [46] consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans.
+This dataset (BMC2021spliceator) consists of “confirmed” error-free splice-site sequences from a diverse set of 148 eukaryotic organisms, including humans.
 The gold standard dataset GS 1 is adopted, which contains an equal number of positive and negative samples, and the F1 score is used as the evaluation metric.
@@
 SpliceAI dataset.
-This dataset [16] also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information.
+This dataset (JAGANATHAN2019SpliceAI) also constructs a binary classification dataset similar to Spliceator, which utilizes the GTEx (Genotype-Tissue Expression) project for RNA sequencing data from various human tissues and the GENCODE V24lift37 canonical annotation for gene structure information.
 SpliceAI also references the ClinVar database to evaluate the clinical significance of predicted splicing variants, which contains information on clinically relevant variants and their associations with diseases.
 This dataset can be regarded as a long-range evaluation and adopts the top-1 AUC-ROC score as the metric.
-Table A6: Full Results of Foundation Models.
-As for DNA and RNA tasks, MCC (%) is reported for Promoter Detection (all), Covid Variants Classification, and Splice Site Reconstructed.
-SRCC (%) is reported for Bacterial and Human Protein Fitness Prediction with DMS.
-Top-1 accuracy (%) is reported for the Central Dogma evaluation.
-The best result is marked as the bold type.
 Method NT-2500M EVO-7B LucaOne Life-Code
@@
 Human Protein 4.7 11.1 19.6 22.4
-Central Dogma -- 75.5 84.8 85.6
+Central Dogma - 75.5 84.8 85.6
+Table A6: Full Results of Foundation Models.
+As for DNA and RNA tasks, MCC (%) is reported for Promoter Detection (all), Covid Variants Classification, and Splice Site Reconstructed.
+SRCC (%) is reported for Bacterial and Human Protein Fitness Prediction with DMS.
+Top-1 accuracy (%) is reported for the Central Dogma evaluation.
+The best result is marked as the bold type.
 B.4 Protein and Multi-omic Tasks
 Zero-shot Protein Fitness Prediction.
-Following EVO [35] and protein language models [27], we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness [38].
+Following EVO (nguyen2024evo) and protein language models (lin2023esm), we employ Deep Mutational Scanning (DMS) studies to evaluate the models’ abilities for protein tasks, which introduce many mutations to a protein coding sequence and then experimentally measure the effects of these mutations (as fitness scores) on various definitions of fitness (icml2022Tranception).
 EVO obtained (DMS) datasets with bacterial (prokaryote) and human (eukaryote) proteins from ProteinGYM at https://proteingym.org.
-To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies [38, 2].
+To adapt this task to nucleotide sequences, EVO proposes to use the wild-type coding sequence and nucleotide mutations reported in the original DMS studies (icml2022Tranception; altae2021widespread).
 For generative pre-trained models such as EVO, we rely on likelihood-based scores under the same masking scheme, assessing how well the model anticipates mutations.
-The model performances of zero-shot function prediction are measured by the strength of Spearman’s Rank Correlation Coefficient (SRCC) that correlates the predicted likelihoods with the experimental fitness measurements.
+The model performances of zero-shot function prediction are measured by the strength of Spearman’s Rank Correlation Coefficient (SRCC), which correlates the predicted likelihoods with the experimental fitness measurements.
 Full results of protein fitness prediction are shown in Table 3 and Table A6, in which our Life-Code achieves balancing performances with bacterial and human proteins and surpasses NT-2500M and EVO-7B.
 ncRNA-Protein Interactions.
-Following LucaOne [13], we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins.
-Since experimentally identifying ncRPI) is typically expensive and time-consuming, the AI-based ncRPI can be a promising task.
+Following LucaOne (he2024lucaone), we consider the multi-omics task of ncRNA-protein interactions (ncRPI), which identifies the interaction strengths between non-coding RNAs (e.g., snRNAs, snoRNAs, miRNAs, and lncRNAs) and proteins.
+Since experimentally identifying ncRPI) It is typically expensive and time-consuming, but the AI-based ncRPI can be a promising task.
 LucaOne proposes a binary classification task involving pairs of ncRNA and Amino Acid sequences (20,824 pairs in total) with top-1 accuracy as the metric.
 Central Dogma Evaluation.
-To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne [13] to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated.
+To evaluate the modeling of the translation rule in the central dogma, we follow LucaOne (he2024lucaone) to conduct a binary classification task with top-1 accuracy, which determines whether the DNA sequences and the given proteins are correlated.
 LucaOne collects a total of 8,533 accurate DNA-protein pairs from 13 species in the NCBI RefSeq database, with each DNA sequence extending to include an additional 100 nucleotides at both the 5’ and 3’ contexts, preserving intron sequences within the data.